### v1

# Thunderscapes: Simulating the Dynamics of Mesoscale Convective System

Abstract
A Mesoscale Convective System (MCS) is a collection of thunderstorms that function as a system, representing a widely discussed phenomenon in both the natural sciences and visual effects industries, and embodying the untamed forces of nature.In this paper, we present the first interactive, physically inspired mesoscale thunderstorms simulation model that integrates Grabowski-style cloud microphysics with atmospheric electrification processes. Our model simulates thunderclouds development and lightning flashes within a unified meteorological framework, providing a realistic and interactive approach for graphical applications. By incorporating key physical principles, it effectively links cloud formation, electrification, and lightning generation. The simulation also encompasses various thunderstorm types and their corresponding lightning activities.
1 INTRODUCTION
Thunderstorms represent the wild power of nature and are a common atmospheric element in visual effects (VFX) industry. Notable works, such as Horizon Forbidden West: Burning Shores and Ghost of Tsushima(see Figure LABEL:fig:thunderstorm_visuals), demonstrate the importance of realistic atmospheric effects. Modern applications demand tools for efficient and realistic simulation of thunderstorms. However, current general purpose VFX software, such as Houdini and Maya, lacks standardized toolkits specifically designed for creating atmospheric phenomena like thunderstorms. Artists often rely on traditional Eulerian fluid dynamics (e.g., stable fluids[Stam(1999)]) to simulate smoke and manually control the evolution of cloud shapes. Lightning, on the other hand, is typically generated using L-system based methods to create fractional geometries. However, these approaches do not account for the natural dynamics and interrelationship between thunderclouds and lightning.
Recent advances in computer graphics research have increasingly focused on cloud dynamics, incorporating atmospheric microphysical processes [Hädrich et al.(2020), Herrera et al.(2021), Amador Herrera et al.(2024)]. However, the critical role of atmospheric electrification in cloud development remains underexplored. This study aims to address this gap by integrating the electrification process, thus enhancing the consistency and realism of phenomena observed during thunderstorm events.
This paper presents an interactive, physically inspired method for simulating thunderstorm development. By coupling cloud microphysics with atmospheric hydrometeor electrification, our approach captures essential processes that contribute to the consistent formation of mesoscale thunderstorm phenomena.
The key contributions include:
  1. (1)
    We present a comprehensive framework for computing thunderstorm microphysics in the atmosphere, incorporating the transport of cloud hydrometeors and their electrification processes.
  2. (2)
    We introduce a lightweight parameterization that enables users to interactively simulate various types of thunderstorms, along with their corresponding lightning effects.
  3. (3)
    We validate our simulation using different meteorological datasets to ensure the generation of physically credible and visually accurate atmospheric effects.
2 RELATED WORK
2.1 Simulating Thunderstorms in Computer Graphics
The simulation of atmospheric phenomena such as thunderstorms has been extensively explored through a variety of computational methods. Webanck et al. [Webanck et al.(2018)] proposed a procedural approach for generating cloudscapes, while Miyazaki et al. [Miyazaki et al.(2002)] simulated cumulus clouds by coupling computational fluid dynamics (CFD) with fundamental water transport equations. Such methods are also employed in Lagrangian-based particle methods. Ferreira et al. [Ferreira Barbosa et al.(2015)] and Zhang et al. [Zhang et al.(2020)] utilized position-based fluids (PBF) for adaptive cloud simulations. Smoothed Particle Hydrodynamics (SPH) techniques, as demonstrated by Goswami and Neyret [Goswami and Neyret(2017)], focus on real-time simulations of convective clouds. Additionally, Vimont et al. [Vimont et al.(2020)] proposed a hybrid, 2D-layered atmospheric model to simulate mesoscale skyscapes, contributing to a more detailed and interactive representation of large-scale cloud structures.
Some studies have specifically focused on the simulation of volcanic cloud dynamics. Lastic et al. [Lastic et al.(2022)] employed Lagrangian dynamics to simulate volcanic plumes and pyroclastic flows, while Pretorius et al. [Pretorius et al.(2024)] integrated volcanic eruptions with atmospheric simulations to produce coherent skyscapes. In their framework, lightning effects were incorporated, but the electrification process, which is essential for lightning formation, was not considered, resulting in a physically inaccurate simulation..
In the domain of lightning simulation, Reed and Wyvill [Reed and Wyvill(1994)], Kim and Lin [Kim and Lin(2007)], and Yun et al. [Yun et al.(2017)] developed methods for the efficient development of lightning branches, contributing to a more realistic representation of the method of electrical discharges.
Recently, more sophisticated microphysical schemes from atmospheric science have been incorporated into computer graphics research. Garcia-Dorado et al. [Garcia-Dorado et al.(2017)] and Hädrich et al. [Hädrich et al.(2020)] adopted the classic Kessler warm cloud microphysics scheme [Kessler(1969)] in cloud simulations, which has been instrumental in capturing the dynamics of cloud formation and precipitation. Herrera et al. [Herrera et al.(2021)] extended cloud simulations to include multiphase cloud dynamics, further enhancing the realism of weather phenomena. Amador Herrera et al. [Amador Herrera et al.(2024)] developed a framework to simulate hurricane and tornado dynamics, incorporating turbulent microphysics to model these extreme atmospheric events with greater accuracy.
2.2 Thunderstorm models in atmospheric sciences
Thunderstorm microphysics and electrification have been widely discussed topics in the field of atmospheric science.
Kessler [Kessler(1969)] proposed a fundamental framework for the distribution and continuity of water substance in atmospheric circulations, which remains influential in the parameterization of the microphysics of warm clouds.
One notable development in cloud microphysics is the work by Grabowski [Grabowski(1998)], who introduced an extended warm cloud microphysics scheme for large-scale tropical circulations. His method divides the parameterization of warm and cold clouds using a temperature interpolation scheme, a significant inspiration for our approach.
In terms of thunderstorm electrification, Solomon et al. [Solomon et al.(2005)] introduced a 1.5-dimensional explicit microphysics thunderstorm model that incorporates a lightning parameterization, addressing key aspects of thunderstorm electrification. Furthermore, Mansell et al. [Mansell et al.(2002)] simulated three-dimensional branched lightning in a numerical thunderstorm model, providing insights on the complex activities of lightning formation. Barthe and Pinty [Barthe and Pinty(2007)] further advanced the field by simulating a supercell storm using a three-dimensional mesoscale model with an explicit lightning flash scheme, capturing the lightning activities specific to supercell thunderstorms.Mansell et al. [Mansell et al.(2010)] also examined the electrification of small thunderstorms using a two-moment bulk microphysics scheme, extending the understanding of lightning activity in small multicell thunderstorms.
Recent studies have focused on exploring the relationship between thundercloud structures and electrification processes. Formenton et al. [Formenton et al.(2013)] used a cloud electrification model to study the relationship between lightning activity and cloud microphysical structure, providing valuable insights into the interaction between cloud formation and lightning generation. More recently, Wu et al. [Wu et al.(2023)] employed coherent Doppler wind lidar to detect and analyze thundercloud structures, offering a novel method for observing thunderstorm activities.
These contributions highlight the ongoing efforts to simulate and understand thunderstorm activities, including cloud microphysics, lightning activity, and the complex interactions between these phenomena.
Our work focuses on the interactive simulation of common thunderstorm types, based on the standard categories in Mesoscale Convective Systems (MCS). According to the National Severe Storms Laboratory (NSSL) [^1] https://www.nssl.noaa.gov/education/svrwx101/thunderstorms/types/, these types include single cell, multi-cell, squall line, and supercell thunderstorms. Additionally, our simulation includes common phenomena generated by thundercloud electrification, such as cloud-to-ground (CG) and intra-cloud (IC) lightning [^2] https://www.nssl.noaa.gov/education/svrwx101/lightning/types/.
Figure 1: Schematic illustration of our thunderstorm microphysics scheme for hydrometeor phase transitions and the electrification process. The details will be further discussed in Chapter 4. (I) Blue frame: the transport of water substances during thundercloud development; (II)Orange frame: the accumulation of charge density due to the collision-coalescence of hydrometeors; (III) Red frame: the static charge, once it exceeds a given electric field threshold, triggers the lightning discharge process; (IV) Blue arrow: the neutralization of hydrometeors following the lightning flash.
3 OVERVIEW
The primary motivation of our approach is to create consistent atmospheric phenomena during thunderstorms, particularly focusing on thundercloud development and dissipation, along with lightning flashes resulting from the thundercloud electrification processes,as illustrated in Figure 1. By simulating both the microphysics of thunderstorms and their electrodynamic properties, we aim to achieve realistic, interactive weather simulations suitable for graphical applications.
Our model integrates key physical principles to simulate thunderstorm behavior at a mesoscale level. The thunderstorm microphysics component models atmospheric dynamics by coupling a Grabowski-style extended warm cloud microphysics scheme with hydrometeor electrification processing. This approach ensures that the microphysical processes driving cloud formation and growth are consistently linked to the electrification necessary for lightning generation. Additionally, we utilize the compact Poisson filter in our fluid dynamics algorithm to create an efficient simulation of the interactions between cloud particles and atmospheric forces. The compact Poisson filter provides an optimized balance between spectral and iterative methods, enabling high-performance simulations while maintaining stability and accuracy, which is essential for interactive graphics applications.
The key atmospheric quantities driving our model include vapor, cloud water, ice, precipitated rain , precipitated snow , static charge (electric charge distribution within the thundercloud, crucial for lightning formation), and lightning (the flash itself, including both its onset and development).
The simulation is capable of modeling various thundercloud types, including single-cell (a simple, isolated thunderstorm cell), multicell (a cluster of thunderstorm cells with varying stages of development), squall line (a collection of storms that form a line hundreds of miles long, often characterized by a shelf cloud feature), supercell (a large, rotating thunderstorm with the potential to produce tornadoes), and atmospheric river (a narrow corridor of concentrated moisture that can fuel intense thunderstorms).
The lightning types simulated in our model include cloud-to-ground lightning (lightning that strikes from the cloud to the ground) and intra-cloud lightning (lightning occurring between different parts of the same cloud).For the simulation of lightning channel growth, we utilize the Dielectric Breakdown Model (DBM)[Kim and Lin(2007)], which captures the electrodynamics involved in the development of lightning.
The validation of our model is performed by comparing simulated results with real-world weather data. This includes cloud fraction profiles to assess cloud formation and structure, temperature, cloud coverage, and relative humidity evolution compared with real-time weather data from national weather services [^3] https://www.visualcrossing.com/weather/weather-data-services/, and lightning flash rate over time, evaluated against observed lightning activity.
Figure 2: Illustration of our thunderstorm microphysics scheme, which integrates the Grabowski-style extended warm cloud microphysics with the hydrometeor electrification process.
4 METHODOLOGY
Our microphysics model, shown in Figure 2, illustrates the interactions among hydrometeor phase transitions and the electrification mechanisms that lead to lightning. Key processes include the condensation of water vapor into droplets and ice crystals, which subsequently form precipitation through autoconversion and accretion. Evaporation recycles hydrometeors back into the vapor phase, while collisions and coalescence facilitate charge separation. Lightning occurs when the electric field strength exceeds a critical threshold, resulting in the discharge and redistribution of accumulated charge. This feedback loop provides a conceptual framework for understanding thunderstorm development within a mesoscale convective system.
4.1 Cloud Microphysics
The fundamental warm-cloud microphysics equations describe the evolution of potential temperature, water vapor, cloud condensate, and precipitation:
$\frac{D\theta}{Dt}=\frac{L_{v}\theta}{c_{p}T}(COND-DIFF)$ (1)
$\frac{Dq_{v}}{Dt}=-COND+DIFF$ (2)
$\frac{Dq_{c}}{Dt}=COND-AUTC-ACCR$ (3)
$\frac{Dq_{p}}{Dt}=AUTC+ACCR-DIFF$ (4)
Here, $\theta$ represents potential temperature, $L_{v}$ is the latent heat of condensation or evaporation, and $c_{p}$ is the specific heat capacity at constant pressure. The variables $q_{v}$, $q_{c}$, and $q_{p}$ denote the mixing ratios of water vapor, cloud condensate, and precipitation, respectively. The terms $COND$, $DIFF$, $AUTC$, and $ACCR$ represent the condensation rate, diffusional growth rate of precipitation, autoconversion rate, and accretion rate, respectively.
We adopt the equilibrium approach proposed by Grabowski, incorporating a temperature-dependent factor $\alpha$ to differentiate warm and cold cloud microphysics. Warm clouds dominate above $0^{\circ}\mathrm{C}$ and cold clouds below $-20^{\circ}\mathrm{C}$, with a linear interpolation for intermediate temperatures:
$q_{vs}=\alpha q_{vw}+(1-\alpha)q_{vi}$ (5)
$q_{w}=\alpha q_{c},\,q_{i}=(1-\alpha)q_{c}$ (6)
$q_{r}=\alpha q_{p},\,q_{s}=(1-\alpha)q_{p}$ (7)
Here, $q_{vs}$ represents the total vapor saturation, combining vapor saturation over water ($q_{vw}$) and ice ($q_{vi}$); $q_{w}$ and $q_{i}$ denote cloud water and ice, expressed as fractions of the cloud-ice mixture ($q_{c}$); $q_{r}$ and $q_{s}$ represent rain and snow, derived as fractions of the rain-snow mixture ($q_{p}$).
The transport equations for rain and snow processes include separate contributions from diffusional growth ($DIFF$), autoconversion ($AUTC$), and accretion ($ACCR$) rates:
$DIFF=DIFF_{r}+DIFF_{s}$ (8)
$AUTC=AUTC_{r}+AUTC_{s}$ (9)
$ACCR=ACCR_{r}+ACCR_{s}$ (10)
The terminal velocity $v_{p}$ of precipitation is modeled as a weighted combination of rain ($v_{r}$) and snow ($v_{s}$) velocities, with typical values $v_{r}=-10\,\mathrm{m/s}$ for rain and $v_{s}=-2\,\mathrm{m/s}$:
$v_{p}=\alpha v_{r}+(1-\alpha)v_{s}$ (11)
The saturation vapor mixing ratio $q_{vs}$ combines contributions from vapor saturation over water and ice, weighted by $\alpha$[Yau and Rogers(1996)]:
$q_{vs}=\frac{380.16}{p}\left[\alpha\exp\left(\frac{17.67\cdot T}{T+243.50}\right)+(1-\alpha)\exp\left(\frac{24.46\cdot T}{T+272.62}\right)\right]$ (12)
Here, $p$ represents the pressure, and $T$ denotes the temperature. The saturation vapor is modeled as an exponential distribution for both liquid water and ice.
The diffusional growth rate $DIFF$ accounts for condensation and evaporation rates, integrating water and ice contributions [Dudhia(1989)],where $\beta_{r},\beta_{s}$ are phase-specific coefficients:
$DIFF=\beta_{r}\cdot\left(\alpha\cdot q_{vs}-q_{v}\right)^{+}+\beta_{s}\cdot\left((1-\alpha)\cdot q_{vs}-q_{v}\right)^{+}$ (13)
The autoconversion rate $AUTC$ describes cloud condensate aggregation into precipitation [Kessler(1995), Lin et al.(1983)]:
$AUTC=\beta_{r}\cdot(\alpha\cdot q_{c}-10^{-3})^{+}+\beta_{s}\cdot e^{0.025T}\cdot((1-\alpha)\cdot q_{c}-10^{-3})^{+}$ (14)
The accretion rate $ACCR$ models the collection of cloud condensate by precipitation particles [Morrison et al.(2015)]:
$ACCR=q_{c}\cdot q_{r}\cdot\left(\beta_{r}\cdot\alpha^{2}+\beta_{s}\cdot(1-\alpha)^{2}\right)$ (15)
4.2 Electrification
Based on the Reynolds-Brook theory of thunderstorm electrification [Latham and Miller(1965)], the fair-weather electric field ($E$) induces opposite charges on precipitation particles. Collisions between particles, influenced by their velocities and directions, result in partial charge neutralization and a residual net charge. Water droplet collisions generally coalesce without significant charge separation, while ice particle interactions are limited by brief contact times and low conductivity. Though water-ice and ice-riming collisions contribute, the inductive process alone is insufficient for thunderstorm-scale electric fields.This theory underpins the following mathematical model for electrification, with charge density ($\rho$) defined as:
$\rho=k\alpha(1-\alpha)\left(q_{c}^{2}v_{c}-q_{p}^{2}v_{p}\right)e^{-10q_{v}}$ (16)
where $k$ is a user-defined constant, $v_{c}$ and $v_{p}$ are their respective velocities of cloud-ice mixture and rain-snow mixture.
The threshold electric field ($E_{tr}$) required for lightning initiation depends on the altitude ($\rho_{A}(h)$) as defined by [Marshall et al.(1995)]:
$E_{tr}=\pm 167\,\rho_{A}(h)\quad\text{where}\quad\rho_{A}(h)=1.208\exp\left(\frac{-h}{8.4}\right)$ (17)
Charge neutralization processes, occurring after a lightning discharge as described by [Barthe and Pinty(2007)], are governed by the following relationship:
$\delta\rho=\begin{cases}\pm\left(|\rho|-\rho_{\text{excess}}\right),&\text{if }|\rho|>\rho_{\text{excess}}\\ 0,&\text{if }|\rho|\leq\rho_{\text{excess}}\end{cases}$ (18)
Here, $\delta\rho$ represents the net charge change, and $\rho_{\text{excess}}$ denotes the threshold for excess charge density. The total neutralized charge density, accounting for the collective contributions of lightning growth points, is expressed as:
$\delta\rho_{\text{neut}}=\frac{1}{N}\left|(2\alpha-1)q_{c}q_{p}\right|\sum_{n}\delta\rho$ (19)
Finally, a suppression factor ($S$) is introduced to modulate lightning activity frequency as thunderstorms dissipate. This factor ensures consistent lightning activity during the lifecycle of thunderstorms:
$S=\frac{1}{1+k\cdot|\delta\rho_{\text{neut}}|}$ (20)
4.3 Atmospheric Background
Our atmospheric background is based on the theory proposed by Hädrich et al. [Hädrich et al.(2020)]. Specifically, we assume that the atmosphere is initially electroneutral, with the charge density, denoted as $\rho$, being zero.
The isentropic exponent $\gamma_{\text{th}}$ for the air-water mixture [Anderson(1990)] is calculated as a weighted average of the vapor-specific exponent $\gamma_{\text{vapor}}$ and the air-specific exponent $\gamma_{\text{air}}$, as shown below:
$\gamma_{\text{th}}=Y_{\text{vapor}}\gamma_{\text{vapor}}+\left(1-Y_{\text{vapor}}\right)\gamma_{\text{air}}$ (21)
where $Y_{\text{vapor}}$ represents the mass fraction of water vapor in the air. The values of $\gamma_{\text{air}}$ and $\gamma_{\text{vapor}}$ are taken as 1.4 and 1.33, respectively, based on standard thermodynamic properties of dry air and water vapor.
The atmospheric temperature profile [Atmosphere(1975)], $T(h)$, is modeled by a piecewise function that accounts for the lapse rate, including the effect of the inversion layer at a height $h_{1}$. The temperature at a given altitude $h$ is expressed as:
$T(h)=\begin{cases}T_{0}+\Gamma_{0}h,&0\leq h\leq h_{1}\\ T_{0}+\Gamma_{0}h_{1}+\Gamma_{1}(h-h_{1}),&h_{1}\leq h\end{cases}$ (22)
where $T_{0}$ is the base temperature at sea level, $\Gamma_{0}$ and $\Gamma_{1}$ represent the lapse rates in the lower and upper layers.
The atmospheric pressure profile, $p(h)$, is derived from the hydrostatic equation [Houze Jr(2014)], considering the effect of gravity and the ideal gas law. It is given by:
$p(h)=p_{0}\left(1-\frac{\Gamma h}{T_{0}}\right)^{\frac{g}{RT_{0}}}$ (23)
where $p_{0}$ is the pressure at sea level, $g$ is the acceleration due to gravity, and $R$ is the specific gas constant.
[…]
The mass fraction of water vapor in the humid air, $Y_{\text{vapor}}$, is related to the mole fraction $X_{\text{vapor}}$ by:
[…]
The temperature of the air in the atmosphere can also be related to pressure changes through the isentropic relation, which governs the temperature $T_{\mathrm{th}}(h)$ at height $h$ in terms of the pressure profile. This is given by:
$T_{\mathrm{th}}(h)=T_{0}\left(\frac{p(h)}{p_{0}}\right)^{\frac{\gamma_{\mathrm{th}}-1}{\gamma_{\mathrm{th}}}}$ (26)
Finally, the buoyancy force, which drives the upward movement of thundercloud, is calculated based on Archimedes’ principle and Newton’s second law. The buoyancy force $B(h)$ at height $h$ is given by:
[…]
4.4 Electrodynamics
The electrodynamic behavior of the atmosphere is governed by Maxwell’s equations[Ma et al.(1998)], which define the relationships between the electric field $\mathbf{E}$, charge density $\rho$, and electric potential $V$. Gauss’s law describes how the divergence of $\mathbf{E}$ is proportional to $\rho$:
$\nabla\cdot\mathbf{E}=\frac{\rho}{\epsilon_{0}}$ (28)
where $\epsilon_{0}$ is the permittivity of free space. The electric field is related to the potential $V$ through:
$\mathbf{E}=-\nabla V$ (29)
indicating that $\mathbf{E}$ points in the direction of decreasing potential. Together, these equations describe the electrostatic field and potential in regions with charge density.
To analyze the electric potential in the atmosphere, we solve the Poisson equation, a fundamental equation derived from Gauss’s law in electrostatics. This equation describes the simplified spatial variation of the potential $V$ due to charge density $\rho$[Kim and Lin(2007)]:
$\nabla^{2}V=-4\pi\rho$ (30)
For simulating lightning discharges, the Dielectric Breakdown Model (DBM) is employed[Niemeyer et al.(1984)]. The probability of a discharge at a specific point $i$ depends on the electric potential $V_{i}$, expressed as:
$p_{i}=\frac{\left(V_{i}\right)^{\eta}}{\sum_{j=1}^{n}\left(V_{j}\right)^{\eta}}$ (31)
Here, $\eta$ is a parameter controlling the spatial concentration of the discharge[Kim and Lin(2007)]. Larger values of $\eta$ bias the discharge toward regions of higher potential, while smaller values allow for more diffusive propagation. The sum in the denominator normalizes the probabilities, ensuring that the total probability across all potential locations is unity. This formulation effectively models lightning channel growth, as discharges preferentially propagate along paths with elevated electric potential.
4.5 Fluid Dynamics
The motion of atmospheric fluids is described by the Navier-Stokes equations, which represent the conservation of momentum in a fluid medium [Wendt(2008)]. These equations capture the effects of inertial forces, pressure gradients, viscosity, and external forces:
$\frac{\partial(\rho\mathbf{u})}{\partial t}+\nabla\cdot(\rho\mathbf{u}\mathbf{u})=-\nabla p+\rho\left(\mu\nabla^{2}\mathbf{u}+\mathbf{b}+\mathbf{f}\right),\quad\nabla\cdot\mathbf{u}=0$ (32)
Here, $\rho$ is the fluid density, $\mathbf{u}$ the velocity field, $p$ the pressure, $\mu$ the dynamic viscosity, $\mathbf{b}$ the body force (e.g., gravity), and $\mathbf{f}$ external forces. The incompressibility condition $\nabla\cdot\mathbf{u}=0$ ensures mass conservation for incompressible flows.
For efficient large-scale atmospheric simulations, we employ the compact Poisson filter method[Rabbani et al.(2022)]. This GPU-friendly solver converts iterative Jacobi steps into a single convolution operation, enabling scalable fluid simulations with controllable accuracy and effective boundary condition handling.
The method solves for a scalar field $\Phi$, which represents the field to be solved in the Poisson equation during the pressure projection.Using the symmetric Poisson kernel $K$, the solution is expressed as:
$K*\Phi=\sum_{i=0}^{r}\mathbf{e}_{i}*_{x}\left(\mathbf{e}_{i}*_{y}\left(\mathbf{e}_{i}*_{z}\Phi\right)\right)$ (33)
Here, $\mathbf{e}_{i}$ are the eigenvectors of $K$, $r$ is the matrix rank, and * denotes convolution along spatial dimensions ($x$, $y$, and $z$).
5 ALGORITHMICS
Algorithm 1 Thunderstorm Development Algorithm.
1:Input: Current MCS state $(\theta,p,\rho,u,q_{v},q_{c},q_{p})$.
2:Output: Updated MCS state.
3:procedure
4:     $\theta,p,\rho\leftarrow$ Update atmospheric background conditions Eqs.(22–23)
5:     $u\leftarrow$ Advect and diffuse velocity field
6:     $b\leftarrow$ Compute thermal buoyancy Eq.(27)
7:     $u\leftarrow u+b+f_{w}+f_{v}$triangleright Apply buoyancy, wind, vorticity confinement
8:     $q_{v},q_{c},q_{p}\leftarrow$ Advect hydrometeor quantities
9:     $u\leftarrow$ Pressure projection using the compact poisson filter Eq.(33)
10:     $q_{v},q_{c},q_{p},\theta,\rho\leftarrow$ Thundercloud microphysics update Eqs.(1–15)
11:     if $S\cdot\rho>E_{tr}$ then $\triangleright$ Electrification process Eq.(16–18)
12:         $V,\rho_{\text{neut}}\leftarrow$ lightning discharge Eqs.(30–31),(19)
13:         $S\cdot\rho\leftarrow$ Apply lightning neutralization Eq.(20)
14:     end if
15:end procedure
The theory discussed in the previous section is implemented as a numerical procedure outlined in Algorithm 1. Figure 1 illustrates the interrelationships governing thunderstorm development. We begin by detailing our procedure for modeling the life cycle of a Mesoscale Convective System (MCS). Subsequently, we describe the dynamics of lightning channel growth, which occurs when the electric field threshold is exceeded. Additionally, we present a lightweight parameter toolset designed to facilitate interactive simulation.
5.1 Mesoscale Convective System Cycle
The input fields for the MCS system, depicted in Figure 3, serve as the foundation for initializing state variables, including potential temperature ($\theta$), pressure ($p$), charge density ($\rho$), velocity field ($u$), and hydrometeor quantities ($q_{v}$, $q_{c}$, $q_{p}$).The solver updates these variables iteratively, producing an output that represents the evolved state of the MCS, incorporating changes driven by thundercloud microphysics and the electrification processes.
Figure 3: Different types of sources used as input for our MCS solver,temperature and vapor field are sampled from perlin noise patterns. (a) Height field derived from satellite data, (b) ground temperature distribution, (c) vapor field distribution.
The procedure begins by updating the atmospheric background conditions,through governing equations (Eqs. 22–23).
The velocity field is then advected and diffused.Thermal buoyancy is computed (Eq. 27) and integrated into the velocity field, alongside contributions from wind and vorticity confinement forces.
Hydrometeor quantities ($q_{v}$, $q_{c}$, $q_{p}$) are transported according to the evolving velocity field through advection. Pressure projection, using the compact Poisson filter (Eq. 33), is applied to ensure stability and enforce flow field divergence constraints. Cloud microphysics are subsequently updated, incorporating processes such as condensation, evaporation, and precipitation, governed by parameterized equations (Eqs. 1–15).
Boundary conditions for the vapor field ($q_{v}$) include periodic boundaries in the horizontal directions and a Dirichlet condition at the top. For temperature ($\theta$), Dirichlet conditions apply in the horizontal and top directions, with a Neumann condition at the bottom.
Electrification is initiated when the charge density, modulated by the suppress factor ($S\cdot\rho$), exceeds the electric field threshold ($E_{tr}$, Eqs. 16–18). This triggers lightning discharges, which are detailed in the next subsection. Following lightning activity, the neutralized charge density ($\rho_{\text{neut}}$, Eqs. 30–31, 19) is updated, redistributing charge density ($S\cdot\rho$, Eq. 20). The suppress factor is recalibrated to reflect the diminished likelihood of subsequent lightning events as the system dissipates.
5.2 Lightning Channel Growth
The growth of the lightning channel begins at a trigger point, which is randomly selected from locations where the electric field exceeds the threshold. The spatial domain for modeling the lightning channel growth is typically represented at a subgrid scale within the thunderstorm model, using a coarser grid size. [Mansell et al.(2002)]This is because the lightning channel is much more sparse compared to the broader thundercloud structure, and it only appears in a limited region of the sky.
Figure 4: Initial charge configurations for different types of lightning. (a) CG Lightning, (b) IC Lightning. Red regions represent areas where the electric potential is equal to 1.
Drawing inspiration from the DBM model proposed by [Kim and Lin(2007)], the lightning channel growth is terminated once it reaches a positive charge object. This approach provides greater flexibility in controlling the final structure of the lightning channel, while also improving computational efficiency.
For cloud-to-ground (CG) lightning, a positive charge configuration is assumed to exist on the ground, with the electric potential set to one. In contrast, for intracloud (IC) lightning, a radial positive charge distribution is established within the boundary of the lightning solving grid,as shown in Figure 4.
The evolution of the lightning channel is modeled by solving the Poisson equation (Eq. 30), using the Gauss-Seidel Successive Overrelaxation (SOR) method. The solution is performed on a grid, and then the results are converted from grid points to geometrical points. To achieve a more natural branching structure, we apply a dithering technique to the points.
Additionally, we employ a custom edge transport method to differentiate point groupings, allowing the model to separately control the widths of return strokes and dart leaders. Finally, this process yields the lightning geometry for the current simulation frame, which is then prepared for rendering.
6 VISUALIZATION
Our tool is implemented as a HDA(Houdini Digital Asset) using Houdini 20.0.625’s microsolver framework with OpenCL for GPU acceleration. The hardware setup includes NVIDIA® GeForce® RTX 3060 Ti GPU,13th Gen Intel® Core™ i5-13600KF processor, and 32GB of RAM.
The simulation results demonstrate various thunderstorm and lightning phenomena, integrated with scenarios inspired by real-world weather events. As shown in Table LABEL:tab:lightweight_params, we summarize the lightweight parameter set used for the scenes presented in this section. All renders are produced using Houdini’s native volumetric rendering engine to ensure high-quality visual outputs.
6.1 Thunderstorms Variation
We simulate four common types of thunderstorms that occur within a mesoscale convective system (MCS), capturing their distinctive characteristics and dynamic behaviors:
  • Single Cell: This is the simplest form of thunderstorm, typically short-lived and consisting of a single updraft and downdraft cycle. It often forms in isolated conditions and usually dissipates within an hour.
  • Multicell: Multicell thunderstorms are composed of multiple convective cells at various stages of their life cycle. These storms exhibit greater longevity and intensity compared to single cells, as new cells continuously form along the gust front of the system.
  • Squall Line: A squall line is a collection of thunderstorms that align into a long, narrow band, often stretching hundreds of miles. This type of storm is typically associated with intense winds, heavy rainfall, and the characteristic ”shelf cloud” feature that marks the leading edge of the gust front.
  • Supercell: The supercell is the most intense and organized type of thunderstorm, characterized by a rotating updraft known as a mesocyclone. These storms are capable of producing severe weather phenomena, including large hail, damaging winds, and tornadoes.
These simulated variations allow us to explore the diverse behaviors and impacts of thunderstorms within an MCS. The visual comparison of these thunderstorm types is presented in Figure LABEL:fig:thundercloud_comparison, highlighting their structural differences and unique features.
6.2 Severe Weather Phenomena
To explore the impact of thunderstorms in real-world scenarios, we reference severe weather events in geographically diverse locations. These events illustrate the variability and intensity of thunderstorms in different environments:
These events serve as the basis for our simulations, showcasing the capability of our model to replicate the diverse and complex phenomena associated with severe weather. The corresponding visualizations are presented in Figure LABEL:fig:weather_simulation.
7 VALIDATION
We follow the evaluation methods in [Shen et al.(2020)] and [Cesana et al.(2019)], tracking altitude-based quantitative data to analyze cloud behavior across atmospheric layers. Results are shown in Figure LABEL:fig:cloud_fraction.
Following the methodology in [Herrera et al.(2021)], we analyze the temporal evolution of four storm types: single cell, multicell, squall line, and supercell. Cloud coverage evolution reveals dynamic distribution changes, while temperature evolution highlights thermal dynamics and energy exchanges.
Following the evaultaion method in [Formenton et al.(2013)],We analyze the lightning flash rate evolution to capture thunderstorm electrical activity. The flash rate rises during storm development and falls to zero as it dissipates, reflecting the storm lifecycle and charge dynamics.
We specifically selected four regions frequently studied in thunderstorms meteorology [Uman(2001), Harris and Carvalho(2018)] to validate our model by comparing its simulation results with real-world weather data. Our focus includes cloud cover dynamics, temperature trends, and relative humidity variations.
Figure LABEL:fig:cloud_real illustrates this comparison across Florida, New Mexico, Japan, and California, showcasing the model’s capability to capture diverse climatic behaviors and atmospheric dynamics.
8 CONCLUSION AND FUTURE WORK
We have developed a comprehensive framework for simulating atmospheric phenomena during thunderstorms, focusing on thundercloud development, dissipation, and lightning generation through cloud electrification processes. This framework integrates a Grabowski-style extended warm cloud microphysics scheme with hydrometeor electrification processes to ensure the consistent coupling of cloud formation and lightning dynamics. Additionally, we incorporated a compact Poisson filter into the fluid dynamics solver to achieve efficient, high-performance simulations with a balance of stability and computational accuracy. The model simulates key atmospheric quantities, including microphysical properties such as vapor, cloud water, ice, and precipitated rain and snow, as well as electrodynamic properties like static charge distribution and lightning flashes. Our framework reproduces various thundercloud types, including single-cell, multicell, squall line, supercell, and atmospheric river formations, and simulates lightning dynamics using the Dielectric Breakdown Model (DBM), capturing both cloud-to-ground and intra-cloud lightning. Validation against real-world weather data demonstrates the model’s accuracy in simulating cloud structure, temperature evolution, cloud coverage, relative humidity, and lightning flash rates, with results benchmarked against national weather services and observed lightning activity.
Future work will focus on expanding the framework to incorporate advanced atmospheric concepts and support a broader range of thunderstorm and lightning phenomena. Enhancements to thunderstorm microphysics, such as more detailed electrification processes and turbulence modeling, will improve the realism of lightning generation. The framework will also be extended to simulate additional thunderstorm types, including Mesoscale Convective Complex (MCC), Mesoscale Convective Vortex (MCV), and derechos, broadening its coverage of mesoscale systems. Furthermore, new lightning types such as anvil crawlers, bolts from the blue, and sheet lightning will be introduced, requiring refined electric field and branching models. Finally, scalability and real-time performance will be improved by leveraging advanced numerical techniques and hardware optimizations, enabling more detailed and interactive simulations.

### v3

@@
 Thunderscapes: Simulating the Dynamics of Mesoscale Convective System
-Abstract
-A Mesoscale Convective System (MCS) is a collection of thunderstorms that function as a system, representing a widely discussed phenomenon in both the natural sciences and visual effects industries, and embodying the untamed forces of nature.In this paper, we present the first interactive, physically inspired mesoscale thunderstorms simulation model that integrates Grabowski-style cloud microphysics with atmospheric electrification processes.
-Our model simulates thunderclouds development and lightning flashes within a unified meteorological framework, providing a realistic and interactive approach for graphical applications.
+Abstract.
+A Mesoscale Convective System (MCS) is a collection of thunderstorms that function as a system, representing a widely discussed phenomenon in both the natural sciences and visual effects industries, and embodying the untamed forces of nature.In this paper, we present the first efficient, physically based mesoscale thunderstorms simulation model that integrates Grabowski-style cloud microphysics with hydrometeor electrification processes.
+Our model simulates thunderclouds development and lightning flashes within a unified meteorological framework, providing a realistic and efficient approach for graphical applications.
 By incorporating key physical principles, it effectively links cloud formation, electrification, and lightning generation.
 The simulation also encompasses various thunderstorm types and their corresponding lightning activities.
-1 INTRODUCTION
+For more details, see the dynamic video: Thunderscapes Video.
+Thunderstorm Modeling and Simulation,Fluid Simulation, Cloud Simulation,Lightning Simulation,Weather Simulation,Atmospheric Microphysics,Atmospheric Electrification
+ccs: Computing methodologies Physical simulation
+1. INTRODUCTION
 Thunderstorms represent the wild power of nature and are a common atmospheric element in visual effects (VFX) industry.
-Notable works, such as Horizon Forbidden West: Burning Shores and Ghost of Tsushima(see Figure LABEL:fig:thunderstorm_visuals), demonstrate the importance of realistic atmospheric effects.
+Notable works, such as Horizon Forbidden West: Burning Shores and Ghost of Tsushima(see Figure 2), demonstrate the importance of realistic atmospheric effects.
 Modern applications demand tools for efficient and realistic simulation of thunderstorms.
 However, current general purpose VFX software, such as Houdini and Maya, lacks standardized toolkits specifically designed for creating atmospheric phenomena like thunderstorms.
-Artists often rely on traditional Eulerian fluid dynamics (e.g., stable fluids[Stam(1999)]) to simulate smoke and manually control the evolution of cloud shapes.
-Lightning, on the other hand, is typically generated using L-system based methods to create fractional geometries.
-However, these approaches do not account for the natural dynamics and interrelationship between thunderclouds and lightning.
-Recent advances in computer graphics research have increasingly focused on cloud dynamics, incorporating atmospheric microphysical processes [Hädrich et al.(2020), Herrera et al.(2021), Amador Herrera et al.(2024)].
+(a) Thunderstorm from *Ghost of Tsushima*. ©2020 Sony Interactive Entertainment LLC.
+[…]
+A scenic image from the video game Ghost of Tsushima showing dramatic landscapes and warrior scenes.
+(b) Thunderstorm from *Horizon Forbidden West: Burning Shores*. ©2023 Sony Interactive Entertainment Europe.
+[…]
+A scene from the video game Horizon Forbidden West: Burning Shores, showcasing futuristic landscapes and action.
+Figure 2. Comparison of atmospheric visual styles between *Ghost of Tsushima* and *Horizon Forbidden West: Burning Shores*.
+Recent advances in computer graphics research have increasingly focused on cloud dynamics, incorporating atmospheric microphysical processes (Hädrich et al., 2020; Herrera et al., 2021; Amador Herrera et al., 2024).
 However, the critical role of atmospheric electrification in cloud development remains underexplored.
 This study aims to address this gap by integrating the electrification process, thus enhancing the consistency and realism of phenomena observed during thunderstorm events.
-This paper presents an interactive, physically inspired method for simulating thunderstorm development.
+This paper presents an efficient, physically based method for simulating thunderstorm development.
 By coupling cloud microphysics with atmospheric hydrometeor electrification, our approach captures essential processes that contribute to the consistent formation of mesoscale thunderstorm phenomena.
@@
   2. (2)
-    We introduce a lightweight parameterization that enables users to interactively simulate various types of thunderstorms, along with their corresponding lightning effects.
+    We introduce a lightweight parameterization that enables users to efficiently simulate various types of thunderstorms, along with their corresponding lightning effects.
   3. (3)
     We validate our simulation using different meteorological datasets to ensure the generation of physically credible and visually accurate atmospheric effects.
-2 RELATED WORK
-2.1 Simulating Thunderstorms in Computer Graphics
+2. RELATED WORK
+2.1. Simulating Thunderstorms in Computer Graphics
 The simulation of atmospheric phenomena such as thunderstorms has been extensively explored through a variety of computational methods.
 Webanck et al.
-[Webanck et al.(2018)] proposed a procedural approach for generating cloudscapes, while Miyazaki et al.
-[Miyazaki et al.(2002)] simulated cumulus clouds by coupling computational fluid dynamics (CFD) with fundamental water transport equations.
-Such methods are also employed in Lagrangian-based particle methods.
+(Webanck et al., 2018) proposed a procedural approach for generating cloudscapes, while Miyazaki et al.
+(Miyazaki et al., 2002) simulated cumulus clouds by coupling computational fluid dynamics (CFD) with fundamental water transport equations.
 Ferreira et al.
-[Ferreira Barbosa et al.(2015)] and Zhang et al.
-[Zhang et al.(2020)] utilized position-based fluids (PBF) for adaptive cloud simulations.
-Smoothed Particle Hydrodynamics (SPH) techniques, as demonstrated by Goswami and Neyret [Goswami and Neyret(2017)], focus on real-time simulations of convective clouds.
+(Ferreira Barbosa et al., 2015) and Zhang et al.
+(Zhang et al., 2020) utilized position-based fluids (PBF) for adaptive cloud simulations.
+Smoothed Particle Hydrodynamics (SPH) techniques, as demonstrated by Goswami and Neyret (Goswami and Neyret, 2017), focus on real-time simulations of convective clouds.
 Additionally, Vimont et al.
-[Vimont et al.(2020)] proposed a hybrid, 2D-layered atmospheric model to simulate mesoscale skyscapes, contributing to a more detailed and interactive representation of large-scale cloud structures.
+(Vimont et al., 2020) proposed a hybrid, 2D-layered atmospheric model to simulate mesoscale skyscapes.
 Some studies have specifically focused on the simulation of volcanic cloud dynamics.
 Lastic et al.
-[Lastic et al.(2022)] employed Lagrangian dynamics to simulate volcanic plumes and pyroclastic flows, while Pretorius et al.
-[Pretorius et al.(2024)] integrated volcanic eruptions with atmospheric simulations to produce coherent skyscapes.
-In their framework, lightning effects were incorporated, but the electrification process, which is essential for lightning formation, was not considered, resulting in a physically inaccurate simulation..
-In the domain of lightning simulation, Reed and Wyvill [Reed and Wyvill(1994)], Kim and Lin [Kim and Lin(2007)], and Yun et al.
-[Yun et al.(2017)] developed methods for the efficient development of lightning branches, contributing to a more realistic representation of the method of electrical discharges.
+(Lastic et al., 2022) employed Lagrangian dynamics to simulate volcanic plumes and pyroclastic flows, while Pretorius et al.
+(Pretorius et al., 2024) integrated volcanic eruptions with atmospheric simulations to produce coherent skyscapes.
+In the domain of lightning simulation, Reed and Wyvill (Reed and Wyvill, 1994), Kim and Lin (Kim and Lin, 2007), and Yun et al.
+(Yun et al., 2017) developed methods for the efficient development of lightning branches, contributing to a more realistic representation of the method of electrical discharges.
 Recently, more sophisticated microphysical schemes from atmospheric science have been incorporated into computer graphics research.
 Garcia-Dorado et al.
-[Garcia-Dorado et al.(2017)] and Hädrich et al.
-[Hädrich et al.(2020)] adopted the classic Kessler warm cloud microphysics scheme [Kessler(1969)] in cloud simulations, which has been instrumental in capturing the dynamics of cloud formation and precipitation.
+(Garcia-Dorado et al., 2017) and Hädrich et al.
+(Hädrich et al., 2020) adopted the classic Kessler warm cloud microphysics scheme (Kessler, 1969) in cloud simulations.
 Herrera et al.
-[Herrera et al.(2021)] extended cloud simulations to include multiphase cloud dynamics, further enhancing the realism of weather phenomena.
+(Herrera et al., 2021) extended cloud simulations to include multiphase cloud dynamics.
 Amador Herrera et al.
-[Amador Herrera et al.(2024)] developed a framework to simulate hurricane and tornado dynamics, incorporating turbulent microphysics to model these extreme atmospheric events with greater accuracy.
-2.2 Thunderstorm models in atmospheric sciences
+(Amador Herrera et al., 2024) developed a framework to simulate hurricane and tornado dynamics.
+2.2. Thunderstorm models in atmospheric sciences
 Thunderstorm microphysics and electrification have been widely discussed topics in the field of atmospheric science.
-Kessler [Kessler(1969)] proposed a fundamental framework for the distribution and continuity of water substance in atmospheric circulations, which remains influential in the parameterization of the microphysics of warm clouds.
-One notable development in cloud microphysics is the work by Grabowski [Grabowski(1998)], who introduced an extended warm cloud microphysics scheme for large-scale tropical circulations.
+Kessler (Kessler, 1969) proposed a fundamental framework for the distribution and continuity of water substance in atmospheric circulations, which remains influential in the parameterization of the microphysics of warm clouds.
+One notable development in cloud microphysics is the work by Grabowski (Grabowski, 1998), who introduced an extended warm cloud microphysics scheme for large-scale tropical circulations.
 His method divides the parameterization of warm and cold clouds using a temperature interpolation scheme, a significant inspiration for our approach.
 In terms of thunderstorm electrification, Solomon et al.
-[Solomon et al.(2005)] introduced a 1.5-dimensional explicit microphysics thunderstorm model that incorporates a lightning parameterization, addressing key aspects of thunderstorm electrification.
+(Solomon et al., 2005) introduced a 1.5-dimensional explicit microphysics thunderstorm model that incorporates a lightning parameterization, addressing key aspects of thunderstorm electrification.
 Furthermore, Mansell et al.
-[Mansell et al.(2002)] simulated three-dimensional branched lightning in a numerical thunderstorm model, providing insights on the complex activities of lightning formation.
-Barthe and Pinty [Barthe and Pinty(2007)] further advanced the field by simulating a supercell storm using a three-dimensional mesoscale model with an explicit lightning flash scheme, capturing the lightning activities specific to supercell thunderstorms.Mansell et al.
-[Mansell et al.(2010)] also examined the electrification of small thunderstorms using a two-moment bulk microphysics scheme, extending the understanding of lightning activity in small multicell thunderstorms.
-Recent studies have focused on exploring the relationship between thundercloud structures and electrification processes.
-Formenton et al.
-[Formenton et al.(2013)] used a cloud electrification model to study the relationship between lightning activity and cloud microphysical structure, providing valuable insights into the interaction between cloud formation and lightning generation.
-More recently, Wu et al.
-[Wu et al.(2023)] employed coherent Doppler wind lidar to detect and analyze thundercloud structures, offering a novel method for observing thunderstorm activities.
-These contributions highlight the ongoing efforts to simulate and understand thunderstorm activities, including cloud microphysics, lightning activity, and the complex interactions between these phenomena.
-Our work focuses on the interactive simulation of common thunderstorm types, based on the standard categories in Mesoscale Convective Systems (MCS).
+(Mansell et al., 2002) simulated three-dimensional branched lightning in a numerical thunderstorm model, providing insights on the complex activities of lightning formation.
+Barthe and Pinty (Barthe and Pinty, 2007) further advanced the field by simulating a supercell storm using a three-dimensional mesoscale model with an explicit lightning flash scheme, capturing the lightning activities specific to supercell thunderstorms.Mansell et al.
+(Mansell et al., 2010) also examined the electrification of small thunderstorms using a two-moment bulk microphysics scheme, extending the understanding of lightning activity in small multicell thunderstorms.
+Our work focuses on the efficient simulation of common thunderstorm types, based on the standard categories in Mesoscale Convective Systems (MCS).
 According to the National Severe Storms Laboratory (NSSL) [^1] https://www.nssl.noaa.gov/education/svrwx101/thunderstorms/types/, these types include single cell, multi-cell, squall line, and supercell thunderstorms.
 Additionally, our simulation includes common phenomena generated by thundercloud electrification, such as cloud-to-ground (CG) and intra-cloud (IC) lightning [^2] https://www.nssl.noaa.gov/education/svrwx101/lightning/types/.
-Figure 1: Schematic illustration of our thunderstorm microphysics scheme for hydrometeor phase transitions and the electrification process.
+Figure 3.
+Schematic illustration of our thunderstorm microphysics scheme for hydrometeor phase transitions and the electrification process.
 The details will be further discussed in Chapter 4.
 (I) Blue frame: the transport of water substances during thundercloud development; (II)Orange frame: the accumulation of charge density due to the collision-coalescence of hydrometeors; (III) Red frame: the static charge, once it exceeds a given electric field threshold, triggers the lightning discharge process; (IV) Blue arrow: the neutralization of hydrometeors following the lightning flash.
-3 OVERVIEW
+[…]
+A diagram illustrating the stages of hydrometeor phase transitions and electrification during thunderstorm development.
+3. OVERVIEW
 The primary motivation of our approach is to create consistent atmospheric phenomena during thunderstorms, particularly focusing on thundercloud development and dissipation, along with lightning flashes resulting from the thundercloud electrification processes,as illustrated in Figure 3.
-By simulating both the microphysics of thunderstorms and their electrodynamic properties, we aim to achieve realistic, interactive weather simulations suitable for graphical applications.
+By simulating both the microphysics of thunderstorms and their electrodynamic properties, we aim to achieve realistic, efficient weather simulations suitable for graphical applications.
 Our model integrates key physical principles to simulate thunderstorm behavior at a mesoscale level.
@@
 This approach ensures that the microphysical processes driving cloud formation and growth are consistently linked to the electrification necessary for lightning generation.
-Additionally, we utilize the compact Poisson filter in our fluid dynamics algorithm to create an efficient simulation of the interactions between cloud particles and atmospheric forces.
-The compact Poisson filter provides an optimized balance between spectral and iterative methods, enabling high-performance simulations while maintaining stability and accuracy, which is essential for interactive graphics applications.
 The key atmospheric quantities driving our model include vapor, cloud water, ice, precipitated rain , precipitated snow , static charge (electric charge distribution within the thundercloud, crucial for lightning formation), and lightning (the flash itself, including both its onset and development).
-The simulation is capable of modeling various thundercloud types, including single-cell (a simple, isolated thunderstorm cell), multicell (a cluster of thunderstorm cells with varying stages of development), squall line (a collection of storms that form a line hundreds of miles long, often characterized by a shelf cloud feature), supercell (a large, rotating thunderstorm with the potential to produce tornadoes), and atmospheric river (a narrow corridor of concentrated moisture that can fuel intense thunderstorms).
-The lightning types simulated in our model include cloud-to-ground lightning (lightning that strikes from the cloud to the ground) and intra-cloud lightning (lightning occurring between different parts of the same cloud).For the simulation of lightning channel growth, we utilize the Dielectric Breakdown Model (DBM)[Kim and Lin(2007)], which captures the electrodynamics involved in the development of lightning.
-The validation of our model is performed by comparing simulated results with real-world weather data.
-This includes cloud fraction profiles to assess cloud formation and structure, temperature, cloud coverage, and relative humidity evolution compared with real-time weather data from national weather services [^3] https://www.visualcrossing.com/weather/weather-data-services/, and lightning flash rate over time, evaluated against observed lightning activity.
-Figure 2: Illustration of our thunderstorm microphysics scheme, which integrates the Grabowski-style extended warm cloud microphysics with the hydrometeor electrification process.
-4 METHODOLOGY
+The validation of our model involves presenting spatially simulated thundercloud structures alongside meteorological characteristics and comparing the temporally simulated results with real-world weather data.
+This includes analyzing cloud fraction profiles to assess cloud formation and structure, tracking the evolution of cloud coverage against real-time data from national weather services [^3] https://www.visualcrossing.com/weather/weather-data-services/, and evaluating the temporal variation in lightning flash rates.
+Figure 4. Illustration of our thunderstorm microphysics scheme, which integrates the Grabowski-style extended warm cloud microphysics with the hydrometeor electrification process.
+[…]
+A schematic representation of the thunderstorm microphysics scheme, coupling the extended warm cloud microphysics with the electrification process of hydrometeors.
+4. METHODOLOGY
 Our microphysics model, shown in Figure 4, illustrates the interactions among hydrometeor phase transitions and the electrification mechanisms that lead to lightning.
@@
 This feedback loop provides a conceptual framework for understanding thunderstorm development within a mesoscale convective system.
-4.1 Cloud Microphysics
+4.1. Cloud Microphysics
 The fundamental warm-cloud microphysics equations describe the evolution of potential temperature, water vapor, cloud condensate, and precipitation:
-$\frac{D\theta}{Dt}=\frac{L_{v}\theta}{c_{p}T}(COND-DIFF)$ (1)
-$\frac{Dq_{v}}{Dt}=-COND+DIFF$ (2)
-$\frac{Dq_{c}}{Dt}=COND-AUTC-ACCR$ (3)
-$\frac{Dq_{p}}{Dt}=AUTC+ACCR-DIFF$ (4)
+(1) $\frac{D\theta}{Dt}=\frac{L_{v}\theta}{c_{p}T}(COND-DIFF)$
+(2) $\frac{Dq_{v}}{Dt}=-COND+DIFF$
+(3) $\frac{Dq_{c}}{Dt}=COND-AUTC-ACCR$
+(4) $\frac{Dq_{p}}{Dt}=AUTC+ACCR-DIFF$
 Here, $\theta$ represents potential temperature, $L_{v}$ is the latent heat of condensation or evaporation, and $c_{p}$ is the specific heat capacity at constant pressure.
@@
 Warm clouds dominate above $0^{\circ}\mathrm{C}$ and cold clouds below $-20^{\circ}\mathrm{C}$, with a linear interpolation for intermediate temperatures:
-$q_{vs}=\alpha q_{vw}+(1-\alpha)q_{vi}$ (5)
-$q_{w}=\alpha q_{c},\,q_{i}=(1-\alpha)q_{c}$ (6)
-$q_{r}=\alpha q_{p},\,q_{s}=(1-\alpha)q_{p}$ (7)
+(5) $q_{vs}=\alpha q_{vw}+(1-\alpha)q_{vi}$
+(6) $q_{w}=\alpha q_{c},\,q_{i}=(1-\alpha)q_{c}$
+(7) $q_{r}=\alpha q_{p},\,q_{s}=(1-\alpha)q_{p}$
 Here, $q_{vs}$ represents the total vapor saturation, combining vapor saturation over water ($q_{vw}$) and ice ($q_{vi}$); $q_{w}$ and $q_{i}$ denote cloud water and ice, expressed as fractions of the cloud-ice mixture ($q_{c}$); $q_{r}$ and $q_{s}$ represent rain and snow, derived as fractions of the rain-snow mixture ($q_{p}$).
 The transport equations for rain and snow processes include separate contributions from diffusional growth ($DIFF$), autoconversion ($AUTC$), and accretion ($ACCR$) rates:
-$DIFF=DIFF_{r}+DIFF_{s}$ (8)
-$AUTC=AUTC_{r}+AUTC_{s}$ (9)
-$ACCR=ACCR_{r}+ACCR_{s}$ (10)
+(8) $DIFF=DIFF_{r}+DIFF_{s}$
+(9) $AUTC=AUTC_{r}+AUTC_{s}$
+(10) $ACCR=ACCR_{r}+ACCR_{s}$
 The terminal velocity $v_{p}$ of precipitation is modeled as a weighted combination of rain ($v_{r}$) and snow ($v_{s}$) velocities, with typical values $v_{r}=-10\,\mathrm{m/s}$ for rain and $v_{s}=-2\,\mathrm{m/s}$:
-$v_{p}=\alpha v_{r}+(1-\alpha)v_{s}$ (11)
-The saturation vapor mixing ratio $q_{vs}$ combines contributions from vapor saturation over water and ice, weighted by $\alpha$[Yau and Rogers(1996)]:
-$q_{vs}=\frac{380.16}{p}\left[\alpha\exp\left(\frac{17.67\cdot T}{T+243.50}\right)+(1-\alpha)\exp\left(\frac{24.46\cdot T}{T+272.62}\right)\right]$ (12)
+(11) $v_{p}=\alpha v_{r}+(1-\alpha)v_{s}$
+The saturation vapor mixing ratio $q_{vs}$ combines contributions from vapor saturation over water and ice, weighted by $\alpha$(Yau and Rogers, 1996):
+(12) $q_{vs}=\frac{380.16}{p}\left[\alpha\exp\left(\frac{17.67\cdot T}{T+243.50}\right)+(1-\alpha)\exp\left(\frac{24.46\cdot T}{T+272.62}\right)\right]$
 Here, $p$ represents the pressure, and $T$ denotes the temperature. The saturation vapor is modeled as an exponential distribution for both liquid water and ice.
-The diffusional growth rate $DIFF$ accounts for condensation and evaporation rates, integrating water and ice contributions [Dudhia(1989)],where $\beta_{r},\beta_{s}$ are phase-specific coefficients:
-$DIFF=\beta_{r}\cdot\left(\alpha\cdot q_{vs}-q_{v}\right)^{+}+\beta_{s}\cdot\left((1-\alpha)\cdot q_{vs}-q_{v}\right)^{+}$ (13)
-The autoconversion rate $AUTC$ describes cloud condensate aggregation into precipitation [Kessler(1995), Lin et al.(1983)]:
-$AUTC=\beta_{r}\cdot(\alpha\cdot q_{c}-10^{-3})^{+}+\beta_{s}\cdot e^{0.025T}\cdot((1-\alpha)\cdot q_{c}-10^{-3})^{+}$ (14)
-The accretion rate $ACCR$ models the collection of cloud condensate by precipitation particles [Morrison et al.(2015)]:
-$ACCR=q_{c}\cdot q_{r}\cdot\left(\beta_{r}\cdot\alpha^{2}+\beta_{s}\cdot(1-\alpha)^{2}\right)$ (15)
-4.2 Electrification
-Based on the Reynolds-Brook theory of thunderstorm electrification [Latham and Miller(1965)], the fair-weather electric field ($E$) induces opposite charges on precipitation particles.
+The diffusional growth rate $DIFF$ accounts for condensation and evaporation rates, integrating water and ice contributions (Dudhia, 1989),where $\beta_{r},\beta_{s}$ are phase-specific coefficients:
+(13) $DIFF=\beta_{r}\cdot\left(\alpha\cdot q_{vs}-q_{v}\right)^{+}+\beta_{s}\cdot\left((1-\alpha)\cdot q_{vs}-q_{v}\right)^{+}$
+The autoconversion rate $AUTC$ describes cloud condensate aggregation into precipitation (Kessler, 1995; Lin et al., 1983):
+(14) $AUTC=\beta_{r}\cdot(\alpha\cdot q_{c}-10^{-3})^{+}+\beta_{s}\cdot e^{0.025T}\cdot((1-\alpha)\cdot q_{c}-10^{-3})^{+}$
+The accretion rate $ACCR$ models the collection of cloud condensate by precipitation particles (Morrison et al., 2015):
+(15) $ACCR=q_{c}\cdot q_{r}\cdot\left(\beta_{r}\cdot\alpha^{2}+\beta_{s}\cdot(1-\alpha)^{2}\right)$
+4.2. Electrification
+Based on the Reynolds-Brook theory of thunderstorm electrification (Latham and Miller, 1965), the fair-weather electric field ($E$) induces opposite charges on precipitation particles.
 Collisions between particles, influenced by their velocities and directions, result in partial charge neutralization and a residual net charge.
-Water droplet collisions generally coalesce without significant charge separation, while ice particle interactions are limited by brief contact times and low conductivity.
-Though water-ice and ice-riming collisions contribute, the inductive process alone is insufficient for thunderstorm-scale electric fields.This theory underpins the following mathematical model for electrification, with charge density ($\rho$) defined as:
-$\rho=k\alpha(1-\alpha)\left(q_{c}^{2}v_{c}-q_{p}^{2}v_{p}\right)e^{-10q_{v}}$ (16)
+This theory underpins the following mathematical model for electrification, with charge density ($\rho$) defined as:
+(16) $\rho=k\alpha(1-\alpha)\left(q_{c}^{2}v_{c}-q_{p}^{2}v_{p}\right)e^{-10q_{v}}$
 where $k$ is a user-defined constant, $v_{c}$ and $v_{p}$ are their respective velocities of cloud-ice mixture and rain-snow mixture.
-The threshold electric field ($E_{tr}$) required for lightning initiation depends on the altitude ($\rho_{A}(h)$) as defined by [Marshall et al.(1995)]:
-$E_{tr}=\pm 167\,\rho_{A}(h)\quad\text{where}\quad\rho_{A}(h)=1.208\exp\left(\frac{-h}{8.4}\right)$ (17)
-Charge neutralization processes, occurring after a lightning discharge as described by [Barthe and Pinty(2007)], are governed by the following relationship:
-$\delta\rho=\begin{cases}\pm\left(|\rho|-\rho_{\text{excess}}\right),&\text{if }|\rho|>\rho_{\text{excess}}\\ 0,&\text{if }|\rho|\leq\rho_{\text{excess}}\end{cases}$ (18)
+The threshold electric field ($E_{tr}$) required for lightning initiation depends on the altitude ($\rho_{A}(h)$) as defined by (Marshall et al., 1995):
+(17) $E_{tr}=\pm 167\,\rho_{A}(h)\quad\text{where}\quad\rho_{A}(h)=1.208\exp\left(\frac{-h}{8.4}\right)$
+Charge neutralization processes, occurring after a lightning discharge as described by (Barthe and Pinty, 2007), are governed by the following relationship:
+(18) $\delta\rho=\begin{cases}\pm\left(|\rho|-\rho_{\text{excess}}\right),&\text{if }|\rho|>\rho_{\text{excess}}\\ 0,&\text{if }|\rho|\leq\rho_{\text{excess}}\end{cases}$
 Here, $\delta\rho$ represents the net charge change, and $\rho_{\text{excess}}$ denotes the threshold for excess charge density.
 The total neutralized charge density, accounting for the collective contributions of lightning growth points, is expressed as:
-$\delta\rho_{\text{neut}}=\frac{1}{N}\left|(2\alpha-1)q_{c}q_{p}\right|\sum_{n}\delta\rho$ (19)
+(19) $\delta\rho_{\text{neut}}=\frac{1}{N}\left|(2\alpha-1)q_{c}q_{p}\right|\sum_{n}\delta\rho$
 Finally, a suppression factor ($S$) is introduced to modulate lightning activity frequency as thunderstorms dissipate.
 This factor ensures consistent lightning activity during the lifecycle of thunderstorms:
-$S=\frac{1}{1+k\cdot|\delta\rho_{\text{neut}}|}$ (20)
-4.3 Atmospheric Background
+(20) $S=\frac{1}{1+k\cdot|\delta\rho_{\text{neut}}|}$
+4.3. Atmospheric Background
 Our atmospheric background is based on the theory proposed by Hädrich et al.
-[Hädrich et al.(2020)].
+(Hädrich et al., 2020).
 Specifically, we assume that the atmosphere is initially electroneutral, with the charge density, denoted as $\rho$, being zero.
-The isentropic exponent $\gamma_{\text{th}}$ for the air-water mixture [Anderson(1990)] is calculated as a weighted average of the vapor-specific exponent $\gamma_{\text{vapor}}$ and the air-specific exponent $\gamma_{\text{air}}$, as shown below:
-$\gamma_{\text{th}}=Y_{\text{vapor}}\gamma_{\text{vapor}}+\left(1-Y_{\text{vapor}}\right)\gamma_{\text{air}}$ (21)
+The isentropic exponent $\gamma_{\text{th}}$ for the air-water mixture (Anderson, 1990) is calculated as a weighted average of the vapor-specific exponent $\gamma_{\text{vapor}}$ and the air-specific exponent $\gamma_{\text{air}}$, as shown below:
+(21) $\gamma_{\text{th}}=Y_{\text{vapor}}\gamma_{\text{vapor}}+\left(1-Y_{\text{vapor}}\right)\gamma_{\text{air}}$
 where $Y_{\text{vapor}}$ represents the mass fraction of water vapor in the air.
 The values of $\gamma_{\text{air}}$ and $\gamma_{\text{vapor}}$ are taken as 1.4 and 1.33, respectively, based on standard thermodynamic properties of dry air and water vapor.
-The atmospheric temperature profile [Atmosphere(1975)], $T(h)$, is modeled by a piecewise function that accounts for the lapse rate, including the effect of the inversion layer at a height $h_{1}$.
+The atmospheric temperature profile (Atmosphere, 1975), $T(h)$, is modeled by a piecewise function that accounts for the lapse rate, including the effect of the inversion layer at a height $h_{1}$.
 The temperature at a given altitude $h$ is expressed as:
-$T(h)=\begin{cases}T_{0}+\Gamma_{0}h,&0\leq h\leq h_{1}\\ T_{0}+\Gamma_{0}h_{1}+\Gamma_{1}(h-h_{1}),&h_{1}\leq h\end{cases}$ (22)
+(22) $T(h)=\begin{cases}T_{0}+\Gamma_{0}h,&0\leq h\leq h_{1}\\ T_{0}+\Gamma_{0}h_{1}+\Gamma_{1}(h-h_{1}),&h_{1}\leq h\end{cases}$
 where $T_{0}$ is the base temperature at sea level, $\Gamma_{0}$ and $\Gamma_{1}$ represent the lapse rates in the lower and upper layers.
-The atmospheric pressure profile, $p(h)$, is derived from the hydrostatic equation [Houze Jr(2014)], considering the effect of gravity and the ideal gas law. It is given by:
-$p(h)=p_{0}\left(1-\frac{\Gamma h}{T_{0}}\right)^{\frac{g}{RT_{0}}}$ (23)
+The atmospheric pressure profile, $p(h)$, is derived from the hydrostatic equation (Houze Jr, 2014), considering the effect of gravity and the ideal gas law. It is given by:
+(23) $p(h)=p_{0}\left(1-\frac{\Gamma h}{T_{0}}\right)^{\frac{g}{RT_{0}}}$
 where $p_{0}$ is the pressure at sea level, $g$ is the acceleration due to gravity, and $R$ is the specific gas constant.
@@
 This is given by:
-$T_{\mathrm{th}}(h)=T_{0}\left(\frac{p(h)}{p_{0}}\right)^{\frac{\gamma_{\mathrm{th}}-1}{\gamma_{\mathrm{th}}}}$ (26)
+(26) $T_{\mathrm{th}}(h)=T_{0}\left(\frac{p(h)}{p_{0}}\right)^{\frac{\gamma_{\mathrm{th}}-1}{\gamma_{\mathrm{th}}}}$
 Finally, the buoyancy force, which drives the upward movement of thundercloud, is calculated based on Archimedes’ principle and Newton’s second law.
@@
 […]
-4.4 Electrodynamics
-The electrodynamic behavior of the atmosphere is governed by Maxwell’s equations[Ma et al.(1998)], which define the relationships between the electric field $\mathbf{E}$, charge density $\rho$, and electric potential $V$.
+4.4. Electrodynamics
+The electrodynamic behavior of the atmosphere is governed by Maxwell’s equations(Ma et al., 1998), which define the relationships between the electric field $\mathbf{E}$, charge density $\rho$, and electric potential $V$.
 Gauss’s law describes how the divergence of $\mathbf{E}$ is proportional to $\rho$:
-$\nabla\cdot\mathbf{E}=\frac{\rho}{\epsilon_{0}}$ (28)
+(28) $\nabla\cdot\mathbf{E}=\frac{\rho}{\epsilon_{0}}$
 where $\epsilon_{0}$ is the permittivity of free space. The electric field is related to the potential $V$ through:
-$\mathbf{E}=-\nabla V$ (29)
+(29) $\mathbf{E}=-\nabla V$
 indicating that $\mathbf{E}$ points in the direction of decreasing potential. Together, these equations describe the electrostatic field and potential in regions with charge density.
 To analyze the electric potential in the atmosphere, we solve the Poisson equation, a fundamental equation derived from Gauss’s law in electrostatics.
-This equation describes the simplified spatial variation of the potential $V$ due to charge density $\rho$[Kim and Lin(2007)]:
-$\nabla^{2}V=-4\pi\rho$ (30)
-For simulating lightning discharges, the Dielectric Breakdown Model (DBM) is employed[Niemeyer et al.(1984)].
+This equation describes the simplified spatial variation of the potential $V$ due to charge density $\rho$(Kim and Lin, 2007):
+(30) $\nabla^{2}V=-4\pi\rho$
+For simulating lightning discharges, the Dielectric Breakdown Model (DBM) is employed(Niemeyer et al., 1984).
 The probability of a discharge at a specific point $i$ depends on the electric potential $V_{i}$, expressed as:
-$p_{i}=\frac{\left(V_{i}\right)^{\eta}}{\sum_{j=1}^{n}\left(V_{j}\right)^{\eta}}$ (31)
-Here, $\eta$ is a parameter controlling the spatial concentration of the discharge[Kim and Lin(2007)].
+(31) $p_{i}=\frac{\left(V_{i}\right)^{\eta}}{\sum_{j=1}^{n}\left(V_{j}\right)^{\eta}}$
+Here, $\eta$ is a parameter controlling the spatial concentration of the discharge(Kim and Lin, 2007).
 Larger values of $\eta$ bias the discharge toward regions of higher potential, while smaller values allow for more diffusive propagation.
-The sum in the denominator normalizes the probabilities, ensuring that the total probability across all potential locations is unity.
-This formulation effectively models lightning channel growth, as discharges preferentially propagate along paths with elevated electric potential.
-4.5 Fluid Dynamics
-The motion of atmospheric fluids is described by the Navier-Stokes equations, which represent the conservation of momentum in a fluid medium [Wendt(2008)].
+4.5. Fluid Dynamics
+The motion of atmospheric fluids is described by the Navier-Stokes equations, which represent the conservation of momentum in a fluid medium (Wendt, 2008).
 These equations capture the effects of inertial forces, pressure gradients, viscosity, and external forces:
-$\frac{\partial(\rho\mathbf{u})}{\partial t}+\nabla\cdot(\rho\mathbf{u}\mathbf{u})=-\nabla p+\rho\left(\mu\nabla^{2}\mathbf{u}+\mathbf{b}+\mathbf{f}\right),\quad\nabla\cdot\mathbf{u}=0$ (32)
+(32) $\frac{\partial(\rho\mathbf{u})}{\partial t}+\nabla\cdot(\rho\mathbf{u}\mathbf{u})=-\nabla p+\rho\left(\mu\nabla^{2}\mathbf{u}+\mathbf{b}+\mathbf{f}\right),\quad\nabla\cdot\mathbf{u}=0$
 Here, $\rho$ is the fluid density, $\mathbf{u}$ the velocity field, $p$ the pressure, $\mu$ the dynamic viscosity, $\mathbf{b}$ the body force (e.g., gravity), and $\mathbf{f}$ external forces.
 The incompressibility condition $\nabla\cdot\mathbf{u}=0$ ensures mass conservation for incompressible flows.
-For efficient large-scale atmospheric simulations, we employ the compact Poisson filter method[Rabbani et al.(2022)].
-This GPU-friendly solver converts iterative Jacobi steps into a single convolution operation, enabling scalable fluid simulations with controllable accuracy and effective boundary condition handling.
-The method solves for a scalar field $\Phi$, which represents the field to be solved in the Poisson equation during the pressure projection.Using the symmetric Poisson kernel $K$, the solution is expressed as:
-$K*\Phi=\sum_{i=0}^{r}\mathbf{e}_{i}*_{x}\left(\mathbf{e}_{i}*_{y}\left(\mathbf{e}_{i}*_{z}\Phi\right)\right)$ (33)
-Here, $\mathbf{e}_{i}$ are the eigenvectors of $K$, $r$ is the matrix rank, and * denotes convolution along spatial dimensions ($x$, $y$, and $z$).
-5 ALGORITHMICS
+5. ALGORITHMICS
 Algorithm 1 Thunderstorm Development Algorithm.
@@
 8:     $q_{v},q_{c},q_{p}\leftarrow$ Advect hydrometeor quantities
-9:     $u\leftarrow$ Pressure projection using the compact poisson filter Eq.(33)
+9:     $u\leftarrow$ Pressure projection
 10:     $q_{v},q_{c},q_{p},\theta,\rho\leftarrow$ Thundercloud microphysics update Eqs.(1–15)
@@
 15:end procedure
-The theory discussed in the previous section is implemented as a numerical procedure outlined in Algorithm 1.
-Figure 1 illustrates the interrelationships governing thunderstorm development.
-We begin by detailing our procedure for modeling the life cycle of a Mesoscale Convective System (MCS).
-Subsequently, we describe the dynamics of lightning channel growth, which occurs when the electric field threshold is exceeded.
-Additionally, we present a lightweight parameter toolset designed to facilitate interactive simulation.
-5.1 Mesoscale Convective System Cycle
+The theoretical framework discussed in the previous section is translated into a numerical procedure, as outlined in Algorithm 1.
+Figure 3 visualizes the key interrelationships governing thunderstorm development.
+We first present our methodology for simulating the life cycle of a Mesoscale Convective System (MCS), followed by an explanation of the dynamics of lightning channel growth triggered when the electric field exceeds a defined threshold.
+5.1. Mesoscale Convective System Cycle
 The input fields for the MCS system, depicted in Figure 5, serve as the foundation for initializing state variables, including potential temperature ($\theta$), pressure ($p$), charge density ($\rho$), velocity field ($u$), and hydrometeor quantities ($q_{v}$, $q_{c}$, $q_{p}$).The solver updates these variables iteratively, producing an output that represents the evolved state of the MCS, incorporating changes driven by thundercloud microphysics and the electrification processes.
-Figure 3: Different types of sources used as input for our MCS solver,temperature and vapor field are sampled from perlin noise patterns.
+Figure 5.
+Different types of sources used as input for our MCS solver,temperature and vapor field are sampled from perlin noise patterns.
 (a) Height field derived from satellite data, (b) ground temperature distribution, (c) vapor field distribution.
@@
 Hydrometeor quantities ($q_{v}$, $q_{c}$, $q_{p}$) are transported according to the evolving velocity field through advection.
-Pressure projection, using the compact Poisson filter (Eq.
-33), is applied to ensure stability and enforce flow field divergence constraints.
+Hydrometeor quantities ($q_{v}$, $q_{c}$, $q_{p}$) are transported through advection based on the evolving velocity field.
+For efficient pressure projection, we employ the compact Poisson filter scheme (Rabbani et al., 2022), a GPU-friendly method for solving large-scale sparse linear systems.
 Cloud microphysics are subsequently updated, incorporating processes such as condensation, evaporation, and precipitation, governed by parameterized equations (Eqs.
 1–15).
-Boundary conditions for the vapor field ($q_{v}$) include periodic boundaries in the horizontal directions and a Dirichlet condition at the top.
-For temperature ($\theta$), Dirichlet conditions apply in the horizontal and top directions, with a Neumann condition at the bottom.
 Electrification is initiated when the charge density, modulated by the suppress factor ($S\cdot\rho$), exceeds the electric field threshold ($E_{tr}$, Eqs.
@@
 The suppress factor is recalibrated to reflect the diminished likelihood of subsequent lightning events as the system dissipates.
-5.2 Lightning Channel Growth
-The growth of the lightning channel begins at a trigger point, which is randomly selected from locations where the electric field exceeds the threshold.
-The spatial domain for modeling the lightning channel growth is typically represented at a subgrid scale within the thunderstorm model, using a coarser grid size.
-[Mansell et al.(2002)]This is because the lightning channel is much more sparse compared to the broader thundercloud structure, and it only appears in a limited region of the sky.
-Figure 4: Initial charge configurations for different types of lightning. (a) CG Lightning, (b) IC Lightning. Red regions represent areas where the electric potential is equal to 1.
-Drawing inspiration from the DBM model proposed by [Kim and Lin(2007)], the lightning channel growth is terminated once it reaches a positive charge object.
-This approach provides greater flexibility in controlling the final structure of the lightning channel, while also improving computational efficiency.
-For cloud-to-ground (CG) lightning, a positive charge configuration is assumed to exist on the ground, with the electric potential set to one.
-In contrast, for intracloud (IC) lightning, a radial positive charge distribution is established within the boundary of the lightning solving grid,as shown in Figure 4.
-The evolution of the lightning channel is modeled by solving the Poisson equation (Eq.
-30), using the Gauss-Seidel Successive Overrelaxation (SOR) method.
-The solution is performed on a grid, and then the results are converted from grid points to geometrical points.
-To achieve a more natural branching structure, we apply a dithering technique to the points.
-Additionally, we employ a custom edge transport method to differentiate point groupings, allowing the model to separately control the widths of return strokes and dart leaders.
-Finally, this process yields the lightning geometry for the current simulation frame, which is then prepared for rendering.
-6 VISUALIZATION
+5.2. Lightning Channel Growth
+The growth of the lightning channel begins at a trigger point, randomly selected from locations where the electric field exceeds a predefined threshold.
+Inspired by the DBM model proposed by (Kim and Lin, 2007), the lightning channel stops growing once it connects to a positively charged object.
+This approach not only enhances flexibility in controlling the final structure of the lightning channel but also improves computational efficiency.
+The positive charge is distributed along the top of the height field.
+The solution is computed on a volumetric grid, and the results are subsequently converted from grid points to geometric points.
+To create a more natural branching structure, a random dithering technique is applied to these points.
+The points are then connected to form the final geometry of the lightning branches.
+6. VISUALIZATION
 Our tool is implemented as a HDA(Houdini Digital Asset) using Houdini 20.0.625’s microsolver framework with OpenCL for GPU acceleration.
-The hardware setup includes NVIDIA® GeForce® RTX 3060 Ti GPU,13th Gen Intel® Core™ i5-13600KF processor, and 32GB of RAM.
+The hardware setup includes NVIDIA® GeForce® RTX A6000 GPU,13th Gen Intel® Core™ i9-13900 processor, and 128GB of RAM.
 The simulation results demonstrate various thunderstorm and lightning phenomena, integrated with scenarios inspired by real-world weather events.
-As shown in Table LABEL:tab:lightweight_params, we summarize the lightweight parameter set used for the scenes presented in this section.
+As shown in Table 1, we summarize the lightweight parameter set used for the scenes presented in this section.
 All renders are produced using Houdini’s native volumetric rendering engine to ensure high-quality visual outputs.
-6.1 Thunderstorms Variation
+Table 1.
+Overview of the lightweight parameter set used in the scenes presented in this section.
+$D$: World dimensions (km); $R$: Volume resolution; $T_{\text{mcs}}$: Computation time in seconds per frame(s).
+Fig. Scene $D$ (km) $R$ $T_{mcs}$ (s)
+6(a) Single Cell $18\times 12\times 18$ $450\times 301\times 450$ 0.48
+6(b) Multicell $20\times 15\times 20$ $500\times 376\times 500$ 0.81
+6(c) Squall Line $30\times 15\times 30$ $500\times 251\times 500$ 0.54
+6(d) Super Cell $30\times 15\times 30$ $500\times 251\times 500$ 0.68
+7(a) Florida $11\times 12\times 11$ $287\times 309\times 287$ 0.24
+7(b) New Mexico $11\times 13\times 11$ $383\times 438\times 383$ 0.59
+7(c) Japan $11\times 13\times 11$ $383\times 444\times 383$ 0.64
+7(d) California $10\times 7\times 10$ $340\times 247\times 340$ 0.28
+6.1. Thunderstorms Variation
 We simulate four common types of thunderstorms that occur within a mesoscale convective system (MCS), capturing their distinctive characteristics and dynamic behaviors:
-  • Single Cell: This is the simplest form of thunderstorm, typically short-lived and consisting of a single updraft and downdraft cycle.
-It often forms in isolated conditions and usually dissipates within an hour.
-  • Multicell: Multicell thunderstorms are composed of multiple convective cells at various stages of their life cycle.
-These storms exhibit greater longevity and intensity compared to single cells, as new cells continuously form along the gust front of the system.
-  • Squall Line: A squall line is a collection of thunderstorms that align into a long, narrow band, often stretching hundreds of miles.
-This type of storm is typically associated with intense winds, heavy rainfall, and the characteristic ”shelf cloud” feature that marks the leading edge of the gust front.
-  • Supercell: The supercell is the most intense and organized type of thunderstorm, characterized by a rotating updraft known as a mesocyclone.
-These storms are capable of producing severe weather phenomena, including large hail, damaging winds, and tornadoes.
+  • Single Cell: Single cell thunderstorms, often referred to as “popcorn” convection, are small, isolated storms characterized by their brief lifespan, typically lasting less than an hour.
+These storms are visually compact with a single updraft and downdraft cycle, forming as isolated towering cumulus clouds.
+  • Multicell: Multicell thunderstorms consist of clusters of individual convective cells, each at varying stages of development.
+They appear as a dynamic structure where new cells continuously form along the gust front, sustaining the system for several hours.
+  • Squall Line: A squall line is a linear arrangement of thunderstorms, visually identifiable by a long and narrow band of cumulonimbus clouds.
+These storms can extend hundreds of miles in length but are typically narrow, often around 10-20 miles wide.
+  • Supercell: The supercell is the most visually striking and organized thunderstorm type, dominated by a large, rotating updraft known as a mesocyclone.
+These storms feature massive, tilted, and rotating cloud structures, often rising up to 50,000 feet.
+The mesocyclone can span up to 10 miles in diameter and persists for hours.
+Refer to caption(a)
+Refer to caption(b)
+Refer to caption(c)
+Refer to caption(d)
+Refer to caption(e)
+Refer to caption(f)
+Refer to caption(g)
+Refer to caption(h)
+Figure 6.
+Comparison between simulation results (a-d) and realistic observations (e-h) for four storm types: single cell, multicell, squall line, and supercell.
+This layout highlights the structural similarities between our simulated storms and their real-world counterparts.
 These simulated variations allow us to explore the diverse behaviors and impacts of thunderstorms within an MCS.
-The visual comparison of these thunderstorm types is presented in Figure LABEL:fig:thundercloud_comparison, highlighting their structural differences and unique features.
-6.2 Severe Weather Phenomena
+The visual comparison of these thunderstorm types is presented in Figure 6, highlighting their structural differences and unique features.
+6.2. Severe Weather Phenomena
 To explore the impact of thunderstorms in real-world scenarios, we reference severe weather events in geographically diverse locations.
 These events illustrate the variability and intensity of thunderstorms in different environments:
+Refer to caption(a)
+Refer to caption(b)
+Refer to caption(c)
+Refer to caption(d)
+Figure 7.
+Simulation cloudscapes inspired by real-world weather events.
+(a) Florida Biscayne National Park, (b) New Mexico Chiricahua Mountains, (c) Japan Fuji-Hakone-Izu National Park, (d) California coastal atmospheric river.
+These visualizations highlight the geographical diversity and meteorological phenomena captured in our simulation framework.
+  • Biscayne National Park, Florida (25.3692, -80.243): On March 13, 1993, a tropical storm occurred in this region during the spring, bringing heavy rainfall and strong winds.
+  • Chiricahua Mountains, New Mexico (35.7943, -106.443): On July 14, 1989, a severe thunderstorm swept through this mountainous area during the summer, producing heavy rain, hail, and strong winds.
+  • Fuji-Hakone-Izu National Park, Japan (37.745, -119.54): On January 10, 2011, a sudden winter ;p‘thunderstorm formed in this region, resulting in brief but intense snowfall.
+  • Monterey Bay, California (36.6197, -121.906): On December 2, 2012, an atmospheric river brought heavy rainfall and thunderstorms along the California coast, accompanied by strong gusty winds.
 These events serve as the basis for our simulations, showcasing the capability of our model to replicate the diverse and complex phenomena associated with severe weather.
-The corresponding visualizations are presented in Figure LABEL:fig:weather_simulation.
-7 VALIDATION
-We follow the evaluation methods in [Shen et al.(2020)] and [Cesana et al.(2019)], tracking altitude-based quantitative data to analyze cloud behavior across atmospheric layers.
-Results are shown in Figure LABEL:fig:cloud_fraction.
-Following the methodology in [Herrera et al.(2021)], we analyze the temporal evolution of four storm types: single cell, multicell, squall line, and supercell.
-Cloud coverage evolution reveals dynamic distribution changes, while temperature evolution highlights thermal dynamics and energy exchanges.
-Following the evaultaion method in [Formenton et al.(2013)],We analyze the lightning flash rate evolution to capture thunderstorm electrical activity.
-The flash rate rises during storm development and falls to zero as it dissipates, reflecting the storm lifecycle and charge dynamics.
-We specifically selected four regions frequently studied in thunderstorms meteorology [Uman(2001), Harris and Carvalho(2018)] to validate our model by comparing its simulation results with real-world weather data.
-Our focus includes cloud cover dynamics, temperature trends, and relative humidity variations.
-Figure LABEL:fig:cloud_real illustrates this comparison across Florida, New Mexico, Japan, and California, showcasing the model’s capability to capture diverse climatic behaviors and atmospheric dynamics.
-8 CONCLUSION AND FUTURE WORK
+The corresponding visualizations are presented in Figure 7.
+7. VALIDATION
+We adopt the evaluation methods outlined in (Shen et al., 2020) and (Cesana et al., 2019), utilizing altitude-based quantitative data to analyze cloud behavior across different atmospheric layers.
+The results are presented in Figure 8, which illustrates the relationship between cloud fraction and storm height.
+Notably, the maximum cloud fraction corresponds to the storm’s peak altitude, consistent with the classical vertical distribution of cumulonimbus clouds in meteorology.
+Figure 8.
+Comparative visualization of cloud fractions under different storm systems.
+From left to right: single-cell, multi-cell, squall line, and supercell.
+Each panel highlights the cloud fraction structure typical of these systems.
+For further validation, we specifically selected four regions frequently studied in thunderstorm meteorology (Uman, 2001; Harris and Carvalho, 2018).
+Following the methodology from (Herrera et al., 2021), we examine the temporal evolution of the thunderstorm described in section 6.2.
+The evolution of cloud coverage is compared with real-world weather data from public sources, showing dynamic changes in the distribution of cloud cover.
+This comparison helps assess the accuracy of the model in capturing cloud coverage patterns observed in nature.we present the maximum storm development height along with the corresponding cloud fraction, providing a detailed overview of the storm’s vertical structure.
+This information is used to validate the model’s ability to simulate the typical altitude and distribution of clouds during storm development.
+In line with the evaluation method in (Formenton et al., 2013), we analyze the evolution of lightning flash rates to capture the electrical activity of thunderstorms.
+The flash rate increases as the storm develops and gradually declines to zero as the storm dissipates.
+This trend reflects the typical lifecycle of a thunderstorm and its charge dynamics.
+Additionally, the frequency of lightning occurrences is examined in relation to altitude.
+Figure 9 provides a comparison across Florida, New Mexico, Japan, and California, illustrating the model’s ability to capture a wide range of climatic behaviors and atmospheric dynamics, and confirming its consistency with observed storm structures and lightning patterns.
+8. CONCLUSION AND FUTURE WORK
 We have developed a comprehensive framework for simulating atmospheric phenomena during thunderstorms, focusing on thundercloud development, dissipation, and lightning generation through cloud electrification processes.
 This framework integrates a Grabowski-style extended warm cloud microphysics scheme with hydrometeor electrification processes to ensure the consistent coupling of cloud formation and lightning dynamics.
-Additionally, we incorporated a compact Poisson filter into the fluid dynamics solver to achieve efficient, high-performance simulations with a balance of stability and computational accuracy.
 The model simulates key atmospheric quantities, including microphysical properties such as vapor, cloud water, ice, and precipitated rain and snow, as well as electrodynamic properties like static charge distribution and lightning flashes.
-Our framework reproduces various thundercloud types, including single-cell, multicell, squall line, supercell, and atmospheric river formations, and simulates lightning dynamics using the Dielectric Breakdown Model (DBM), capturing both cloud-to-ground and intra-cloud lightning.
+Our framework reproduces various thundercloud types, including single-cell, multicell, squall line, supercell, and atmospheric river formations, and simulates lightning dynamics using the Dielectric Breakdown Model (DBM), capturing both CG and IC lightning.
 Validation against real-world weather data demonstrates the model’s accuracy in simulating cloud structure, temperature evolution, cloud coverage, relative humidity, and lightning flash rates, with results benchmarked against national weather services and observed lightning activity.
+Figure 9.
+Comparison of cloud coverage evolution (top row) ,cloud fraction (middle row) and lightning activity evolution(bottom row)across four distinct regions: Florida, New Mexico, Japan, and California.
+The figures showcase the distinct patterns of atmospheric dynamics influenced by regional environmental conditions.
 Future work will focus on expanding the framework to incorporate advanced atmospheric concepts and support a broader range of thunderstorm and lightning phenomena.
@@
 Furthermore, new lightning types such as anvil crawlers, bolts from the blue, and sheet lightning will be introduced, requiring refined electric field and branching models.
-Finally, scalability and real-time performance will be improved by leveraging advanced numerical techniques and hardware optimizations, enabling more detailed and interactive simulations.
+Finally, scalability and real-time performance will be improved by leveraging advanced numerical techniques and hardware optimizations, enabling more detailed and efficient simulations.

### v4

@@
 Abstract.
-A Mesoscale Convective System (MCS) is a collection of thunderstorms that function as a system, representing a widely discussed phenomenon in both the natural sciences and visual effects industries, and embodying the untamed forces of nature.In this paper, we present the first efficient, physically based mesoscale thunderstorms simulation model that integrates Grabowski-style cloud microphysics with hydrometeor electrification processes.
-Our model simulates thunderclouds development and lightning flashes within a unified meteorological framework, providing a realistic and efficient approach for graphical applications.
-By incorporating key physical principles, it effectively links cloud formation, electrification, and lightning generation.
-The simulation also encompasses various thunderstorm types and their corresponding lightning activities.
-For more details, see the dynamic video: Thunderscapes Video.
+A Mesoscale Convective System (MCS) is a collection of thunderstorms operating as a unified system, showcasing nature’s untamed power.
+They represent a phenomenon widely referenced in both the natural sciences and the visual effects (VFX) industries.However, in computer graphics, visually accurate simulation of MCS dynamics remains a significant challenge due to the inherent complexity of atmospheric microphysical processes.To achieve a high level of visual quality while ensuring practical performance, we introduce Thunderscapes, the first physically based simulation framework for visually realistic MCS tailored to graphical applications.Our model integrates mesoscale cloud microphysics with hydrometeor electrification processes to simulate thunderstorm development and lightning flashes.
+By capturing various thunderstorm types and their associated lightning activities, Thunderscapes demonstrates the versatility and physical accuracy of the proposed approach.
 Thunderstorm Modeling and Simulation,Fluid Simulation, Cloud Simulation,Lightning Simulation,Weather Simulation,Atmospheric Microphysics,Atmospheric Electrification
-ccs: Computing methodologies Physical simulation
+copyright: acmlicensedjournalyear: 2025doi: 10.1145/3728312journal: PACMCGITjournalvolume: 8journalnumber: 1article: 7publicationmonth: 5ccs: Computing methodologies Physical simulation
 1. INTRODUCTION
-Thunderstorms represent the wild power of nature and are a common atmospheric element in visual effects (VFX) industry.
-Notable works, such as Horizon Forbidden West: Burning Shores and Ghost of Tsushima(see Figure 2), demonstrate the importance of realistic atmospheric effects.
+Thunderstorms represent the wild power of nature and are a common atmospheric element in the visual effects (VFX) industry.
+Notable works, such as Horizon Forbidden West: Burning Shores and Ghost of Tsushima, demonstrate the importance of realistic atmospheric effects.
 Modern applications demand tools for efficient and realistic simulation of thunderstorms.
-However, current general purpose VFX software, such as Houdini and Maya, lacks standardized toolkits specifically designed for creating atmospheric phenomena like thunderstorms.
-(a) Thunderstorm from *Ghost of Tsushima*. ©2020 Sony Interactive Entertainment LLC.
-[…]
-A scenic image from the video game Ghost of Tsushima showing dramatic landscapes and warrior scenes.
-(b) Thunderstorm from *Horizon Forbidden West: Burning Shores*. ©2023 Sony Interactive Entertainment Europe.
-[…]
-A scene from the video game Horizon Forbidden West: Burning Shores, showcasing futuristic landscapes and action.
-Figure 2. Comparison of atmospheric visual styles between *Ghost of Tsushima* and *Horizon Forbidden West: Burning Shores*.
-Recent advances in computer graphics research have increasingly focused on cloud dynamics, incorporating atmospheric microphysical processes (Hädrich et al., 2020; Herrera et al., 2021; Amador Herrera et al., 2024).
-However, the critical role of atmospheric electrification in cloud development remains underexplored.
-This study aims to address this gap by integrating the electrification process, thus enhancing the consistency and realism of phenomena observed during thunderstorm events.
-This paper presents an efficient, physically based method for simulating thunderstorm development.
-By coupling cloud microphysics with atmospheric hydrometeor electrification, our approach captures essential processes that contribute to the consistent formation of mesoscale thunderstorm phenomena.
-The key contributions include:
+However, current general-purpose VFX software, such as Houdini and Maya, relies on traditional Eulerian fluid dynamics (e.g., stable fluids (Stam, 1999)) to simulate phenomena like smoke or explosions.
+They fall short in capturing the inherent complexity of atmospheric microphysical processes, which are essential for accurately simulating the dynamics of complex thunderstorm systems, such as MCS, and for producing consistent, high-quality results.
+This paper introduces Thunderscapes, a physically based simulation framework for visually realistic MCS.
+By incorporating a Grabowski-style cloud microphysics model (Grabowski, 1998) alongside hydrometeor electrification processes, our approach delivers high visual fidelity while maintaining practical performance levels, capturing the core dynamics of thunderstorm evolution and lightning activities within an MCS.
+Furthermore, the framework integrates seamlessly with Houdini, a widely used tool in the VFX industry, through intuitive and lightweight parameters, allowing artists to utilize its advanced simulation features without requiring expertise in meteorology or knowledge of supercomputing.
+The key contributions of this work are as follows:
   1. (1)
-    We present a comprehensive framework for computing thunderstorm microphysics in the atmosphere, incorporating the transport of cloud hydrometeors and their electrification processes.
+    We develop a realistic, physically based simulation framework for MCS, integrating mesoscale cloud microphysics with hydrometeor electrification processes to simulate diverse thunderstorm phenomena.
   2. (2)
-    We introduce a lightweight parameterization that enables users to efficiently simulate various types of thunderstorms, along with their corresponding lightning effects.
+    We present a lightweight set of parameters that allows artists to intuitively create thunderstorm animations within general-purpose VFX software.
   3. (3)
-    We validate our simulation using different meteorological datasets to ensure the generation of physically credible and visually accurate atmospheric effects.
+    We validate the framework using diverse meteorological datasets, demonstrating its capability to generate visually realistic and physically consistent atmospheric effects.
 2. RELATED WORK
-2.1. Simulating Thunderstorms in Computer Graphics
-The simulation of atmospheric phenomena such as thunderstorms has been extensively explored through a variety of computational methods.
+2.1. Simulating Thunderstorm in Computer Graphics
+The visual simulation of atmospheric phenomena such as convective clouds has been extensively explored through a variety of computational methods.
 Webanck et al.
@@
 (Vimont et al., 2020) proposed a hybrid, 2D-layered atmospheric model to simulate mesoscale skyscapes.
-Some studies have specifically focused on the simulation of volcanic cloud dynamics.
-Lastic et al.
-(Lastic et al., 2022) employed Lagrangian dynamics to simulate volcanic plumes and pyroclastic flows, while Pretorius et al.
-(Pretorius et al., 2024) integrated volcanic eruptions with atmospheric simulations to produce coherent skyscapes.
+However, these approaches lack a detailed focus on cloud microphysics, limiting their ability to produce realistic and complex cloudscapes.
 In the domain of lightning simulation, Reed and Wyvill (Reed and Wyvill, 1994), Kim and Lin (Kim and Lin, 2007), and Yun et al.
@@
 Amador Herrera et al.
-(Amador Herrera et al., 2024) developed a framework to simulate hurricane and tornado dynamics.
+(Amador Herrera et al., 2024) developed a framework to simulate hurricane and tornado dynamics.However, these methods overlook the hydrometeor electrification process, which limits their ability to simulate consistent and realistic thunderstorms with associated lightning phenomena.
 2.2. Thunderstorm models in atmospheric sciences
-Thunderstorm microphysics and electrification have been widely discussed topics in the field of atmospheric science.
-Kessler (Kessler, 1969) proposed a fundamental framework for the distribution and continuity of water substance in atmospheric circulations, which remains influential in the parameterization of the microphysics of warm clouds.
+Thunderstorm microphysics and electrification have been widely discussed topics in the field of atmospheric science.Kessler (Kessler, 1969) proposed a fundamental framework for the distribution and continuity of water substance in atmospheric circulations, which remains influential in the parameterization of the microphysics of warm clouds.
 One notable development in cloud microphysics is the work by Grabowski (Grabowski, 1998), who introduced an extended warm cloud microphysics scheme for large-scale tropical circulations.
@@
 (Mansell et al., 2010) also examined the electrification of small thunderstorms using a two-moment bulk microphysics scheme, extending the understanding of lightning activity in small multicell thunderstorms.
-Our work focuses on the efficient simulation of common thunderstorm types, based on the standard categories in Mesoscale Convective Systems (MCS).
-According to the National Severe Storms Laboratory (NSSL) [^1] https://www.nssl.noaa.gov/education/svrwx101/thunderstorms/types/, these types include single cell, multi-cell, squall line, and supercell thunderstorms.
-Additionally, our simulation includes common phenomena generated by thundercloud electrification, such as cloud-to-ground (CG) and intra-cloud (IC) lightning [^2] https://www.nssl.noaa.gov/education/svrwx101/lightning/types/.
+3. OVERVIEW
+The primary motivation for our approach is to visually capture the realistic development of MCS.
+We propose a physically based simulation framework that is intuitive for artists, enabling the creation of visually realistic thunderstorm animations while managing physical complexity and ensuring practical performance.
+Simulating MCS is challenging due to the complex microphysical and electrostatic dynamics and the interactions among hydrometeors (ice, snow, and rain) that induce charge imbalances.
+Traditional meteorological models capture these phenomena with high accuracy but rely on numerous variables and detailed phase transitions(Barthe and Pinty, 2007), making them computationally expensive and difficult for non-experts to manage.For example,the conventional warm-cloud model typically involves three or four phase transitions and works well for ordinary cumulus clouds, whereas modeling complex thunderstorms generally requires a combined warm-cold cloud scheme with more than ten phase transitions (e.g., melting and riming (Rutledge and Hobbs, 1984)).
+Although such detailed modeling improves numerical weather prediction, its prohibitively long computation times necessitate the use of supercomputers.
+In contrast,our Grabowski-style framework extends the basic warm-cloud model with a temperature interpolation scheme to incorporate cold-cloud processes, capturing the essential material transformations during thunderstorm development with fewer phase transitions.
+This simplification sacrifices some microscopic detail yet preserves a physically accurate macroscopic evolution, reduces the number of user-controlled variables, and ensures practical performance.
+Building on these advantages, we develop a comprehensive framework to describe the essential microphysical processes of hydrometeor phase transitions and electrification.Our model employs six key parameters to characterize the atmospheric state: the vapor field $q_{v}$, cloud field $q_{c}$, precipitation field $q_{p}$, temperature field $\theta$, charge density field $\rho$, and velocity field $\mathbf{u}$.Our simulation framework, as illustrated in Figure 2, consists of the following components:(I) Cloud microphysics, describing hydrometeor phase transitions such as condensation, evaporation, and precipitation; (II) Hydrometeor electrification, modeling charge accumulation during hydrometeor interactions; (III) Lightning discharge, computing charge neutralization as lightning channels propagate; (IV) Atmospheric background, accounting for buoyancy forces arising from temperature gradients; (V) Fluid dynamics, driving field evolution with an Eulerian solver for MCS development.
 Figure 2.
-Schematic illustration of our thunderstorm microphysics scheme for hydrometeor phase transitions and the electrification process.
-The details will be further discussed in Chapter 4.
-(I) Blue frame: the transport of water substances during thundercloud development; (II)Orange frame: the accumulation of charge density due to the collision-coalescence of hydrometeors; (III) Red frame: the static charge, once it exceeds a given electric field threshold, triggers the lightning discharge process; (IV) Blue arrow: the neutralization of hydrometeors following the lightning flash.
+Schematic illustration of our MCS life cycle, encompassing the interplay and feedback among key modules.
+Detailed explanations are provided in Section 4.(I) Cloud Microphysics: Describes the phase transitions of hydrometeors during MCS development.(II) Hydrometeor Electrification: Explains the accumulation of charge density resulting from the collision and coalescence of hydrometeors.(III) Lightning Discharge: Highlights the process of lightning channel formation and propagation, triggered when static charge exceeds a predefined electric field threshold, leading to the neutralization of hydrometeors along the lightning channel.(IV) Atmospheric Background: Introduces vertical buoyancy forces to the system driven by temperature variations caused by hydrometeor phase transitions.(V) Fluid Dynamics: Drives the system’s temporal evolution and introduces forces influencing the dynamics of the MCS life cycle.
 […]
 A diagram illustrating the stages of hydrometeor phase transitions and electrification during thunderstorm development.
-3. OVERVIEW
-The primary motivation of our approach is to create consistent atmospheric phenomena during thunderstorms, particularly focusing on thundercloud development and dissipation, along with lightning flashes resulting from the thundercloud electrification processes,as illustrated in Figure 3.
-By simulating both the microphysics of thunderstorms and their electrodynamic properties, we aim to achieve realistic, efficient weather simulations suitable for graphical applications.
-Our model integrates key physical principles to simulate thunderstorm behavior at a mesoscale level.
-The thunderstorm microphysics component models atmospheric dynamics by coupling a Grabowski-style extended warm cloud microphysics scheme with hydrometeor electrification processing.
-This approach ensures that the microphysical processes driving cloud formation and growth are consistently linked to the electrification necessary for lightning generation.
-The key atmospheric quantities driving our model include vapor, cloud water, ice, precipitated rain , precipitated snow , static charge (electric charge distribution within the thundercloud, crucial for lightning formation), and lightning (the flash itself, including both its onset and development).
-The validation of our model involves presenting spatially simulated thundercloud structures alongside meteorological characteristics and comparing the temporally simulated results with real-world weather data.
-This includes analyzing cloud fraction profiles to assess cloud formation and structure, tracking the evolution of cloud coverage against real-time data from national weather services [^3] https://www.visualcrossing.com/weather/weather-data-services/, and evaluating the temporal variation in lightning flash rates.
-Figure 4. Illustration of our thunderstorm microphysics scheme, which integrates the Grabowski-style extended warm cloud microphysics with the hydrometeor electrification process.
-[…]
-A schematic representation of the thunderstorm microphysics scheme, coupling the extended warm cloud microphysics with the electrification process of hydrometeors.
 4. METHODOLOGY
-Our microphysics model, shown in Figure 4, illustrates the interactions among hydrometeor phase transitions and the electrification mechanisms that lead to lightning.
+Our microphysics model, depicted in Figure 2, illustrates the interactions among hydrometeor phase transitions and electrification mechanisms.
 Key processes include the condensation of water vapor into droplets and ice crystals, which subsequently form precipitation through autoconversion and accretion.
-Evaporation recycles hydrometeors back into the vapor phase, while collisions and coalescence facilitate charge separation.
-Lightning occurs when the electric field strength exceeds a critical threshold, resulting in the discharge and redistribution of accumulated charge.
-This feedback loop provides a conceptual framework for understanding thunderstorm development within a mesoscale convective system.
+Evaporation recycles hydrometeors back into the vapor phase, as detailed in Section 4.1.
+Collisions and coalescence facilitate charge separation, leading to lightning when the electric field strength exceeds a critical threshold, as described in Section 4.2.
+The buoyancy forces generated by atmospheric temperature variations are discussed in Section 4.3, while the system’s dynamics updates are outlined in Section 4.4.
+This feedback loop provides a comprehensive framework for understanding thunderstorm development within an MCS.
 4.1. Cloud Microphysics
-The fundamental warm-cloud microphysics equations describe the evolution of potential temperature, water vapor, cloud condensate, and precipitation:
-(1) $\frac{D\theta}{Dt}=\frac{L_{v}\theta}{c_{p}T}(COND-DIFF)$
-(2) $\frac{Dq_{v}}{Dt}=-COND+DIFF$
-(3) $\frac{Dq_{c}}{Dt}=COND-AUTC-ACCR$
-(4) $\frac{Dq_{p}}{Dt}=AUTC+ACCR-DIFF$
-Here, $\theta$ represents potential temperature, $L_{v}$ is the latent heat of condensation or evaporation, and $c_{p}$ is the specific heat capacity at constant pressure.
-The variables $q_{v}$, $q_{c}$, and $q_{p}$ denote the mixing ratios of water vapor, cloud condensate, and precipitation, respectively.
-The terms $COND$, $DIFF$, $AUTC$, and $ACCR$ represent the condensation rate, diffusional growth rate of precipitation, autoconversion rate, and accretion rate, respectively.
-We adopt the equilibrium approach proposed by Grabowski, incorporating a temperature-dependent factor $\alpha$ to differentiate warm and cold cloud microphysics.
-Warm clouds dominate above $0^{\circ}\mathrm{C}$ and cold clouds below $-20^{\circ}\mathrm{C}$, with a linear interpolation for intermediate temperatures:
-(5) $q_{vs}=\alpha q_{vw}+(1-\alpha)q_{vi}$
-(6) $q_{w}=\alpha q_{c},\,q_{i}=(1-\alpha)q_{c}$
-(7) $q_{r}=\alpha q_{p},\,q_{s}=(1-\alpha)q_{p}$
-Here, $q_{vs}$ represents the total vapor saturation, combining vapor saturation over water ($q_{vw}$) and ice ($q_{vi}$); $q_{w}$ and $q_{i}$ denote cloud water and ice, expressed as fractions of the cloud-ice mixture ($q_{c}$); $q_{r}$ and $q_{s}$ represent rain and snow, derived as fractions of the rain-snow mixture ($q_{p}$).
-The transport equations for rain and snow processes include separate contributions from diffusional growth ($DIFF$), autoconversion ($AUTC$), and accretion ($ACCR$) rates:
-(8) $DIFF=DIFF_{r}+DIFF_{s}$
-(9) $AUTC=AUTC_{r}+AUTC_{s}$
-(10) $ACCR=ACCR_{r}+ACCR_{s}$
-The terminal velocity $v_{p}$ of precipitation is modeled as a weighted combination of rain ($v_{r}$) and snow ($v_{s}$) velocities, with typical values $v_{r}=-10\,\mathrm{m/s}$ for rain and $v_{s}=-2\,\mathrm{m/s}$:
-(11) $v_{p}=\alpha v_{r}+(1-\alpha)v_{s}$
-The saturation vapor mixing ratio $q_{vs}$ combines contributions from vapor saturation over water and ice, weighted by $\alpha$(Yau and Rogers, 1996):
-(12) $q_{vs}=\frac{380.16}{p}\left[\alpha\exp\left(\frac{17.67\cdot T}{T+243.50}\right)+(1-\alpha)\exp\left(\frac{24.46\cdot T}{T+272.62}\right)\right]$
-Here, $p$ represents the pressure, and $T$ denotes the temperature. The saturation vapor is modeled as an exponential distribution for both liquid water and ice.
-The diffusional growth rate $DIFF$ accounts for condensation and evaporation rates, integrating water and ice contributions (Dudhia, 1989),where $\beta_{r},\beta_{s}$ are phase-specific coefficients:
-(13) $DIFF=\beta_{r}\cdot\left(\alpha\cdot q_{vs}-q_{v}\right)^{+}+\beta_{s}\cdot\left((1-\alpha)\cdot q_{vs}-q_{v}\right)^{+}$
-The autoconversion rate $AUTC$ describes cloud condensate aggregation into precipitation (Kessler, 1995; Lin et al., 1983):
-(14) $AUTC=\beta_{r}\cdot(\alpha\cdot q_{c}-10^{-3})^{+}+\beta_{s}\cdot e^{0.025T}\cdot((1-\alpha)\cdot q_{c}-10^{-3})^{+}$
-The accretion rate $ACCR$ models the collection of cloud condensate by precipitation particles (Morrison et al., 2015):
-(15) $ACCR=q_{c}\cdot q_{r}\cdot\left(\beta_{r}\cdot\alpha^{2}+\beta_{s}\cdot(1-\alpha)^{2}\right)$
-4.2. Electrification
-Based on the Reynolds-Brook theory of thunderstorm electrification (Latham and Miller, 1965), the fair-weather electric field ($E$) induces opposite charges on precipitation particles.
-Collisions between particles, influenced by their velocities and directions, result in partial charge neutralization and a residual net charge.
-This theory underpins the following mathematical model for electrification, with charge density ($\rho$) defined as:
-(16) $\rho=k\alpha(1-\alpha)\left(q_{c}^{2}v_{c}-q_{p}^{2}v_{p}\right)e^{-10q_{v}}$
-where $k$ is a user-defined constant, $v_{c}$ and $v_{p}$ are their respective velocities of cloud-ice mixture and rain-snow mixture.
-The threshold electric field ($E_{tr}$) required for lightning initiation depends on the altitude ($\rho_{A}(h)$) as defined by (Marshall et al., 1995):
-(17) $E_{tr}=\pm 167\,\rho_{A}(h)\quad\text{where}\quad\rho_{A}(h)=1.208\exp\left(\frac{-h}{8.4}\right)$
-Charge neutralization processes, occurring after a lightning discharge as described by (Barthe and Pinty, 2007), are governed by the following relationship:
-(18) $\delta\rho=\begin{cases}\pm\left(|\rho|-\rho_{\text{excess}}\right),&\text{if }|\rho|>\rho_{\text{excess}}\\ 0,&\text{if }|\rho|\leq\rho_{\text{excess}}\end{cases}$
-Here, $\delta\rho$ represents the net charge change, and $\rho_{\text{excess}}$ denotes the threshold for excess charge density.
+The fundamental warm-cloud microphysics equations govern the evolution of potential temperature, water vapor, cloud condensate, and precipitation.
+These equations are formulated using the material derivative (Kundu et al., 2012):
+(1) $\frac{\mathrm{D}\phi}{\mathrm{D}t}=\frac{\partial\phi}{\partial t}+\boldsymbol{u}\cdot\nabla\phi,$
+which provides the foundation for expressing the microphysics equations:
+(2) $\frac{\mathrm{D}\theta}{\mathrm{D}t}=\frac{L_{v}\theta}{C_{p}T}(C_{C}-D_{P}),$
+(3) $\frac{\mathrm{D}q_{v}}{\mathrm{D}t}=-C_{C}+D_{P},$
+(4) $\frac{\mathrm{D}q_{c}}{\mathrm{D}t}=C_{C}-A_{C}-G_{C},$
+(5) $\frac{\mathrm{D}q_{p}}{\mathrm{D}t}=A_{C}+G_{C}-D_{P},$
+where $\theta$ represents the potential temperature, $L_{V}$ is the latent heat of condensation or evaporation, $T$ is the absolute temperature, and $C_{p}$ is the specific heat capacity at constant pressure.
+Additionally, $q_{v}$, $q_{c}$, and $q_{p}$ denote the mixing ratios of water vapor, cloud condensate, and precipitation, respectively.
+The terms $C_{C}$, $D_{P}$, $A_{C}$, and $G_{C}$ correspond to the condensation rate, diffusional growth rate of precipitation, autoconversion rate, and accretion rate.
+We adopt the equilibrium approach proposed by Grabowski (Grabowski, 1998), incorporating a temperature-dependent factor $\alpha$ to differentiate warm and cold cloud microphysics.
+The parameter $\alpha$ varies linearly with temperature: warm clouds dominate above $0^{\circ}\mathrm{C}$, cold clouds dominate below $-20^{\circ}\mathrm{C}$.
+The unified relationship for vapor saturation, cloud, and precipitation components is expressed as:
+(6) $q_{mc}=\alpha q_{wc}+(1-\alpha)q_{cc},$
+where $q_{mc}$ represents the total quantity of a variable $q_{vs},q_{c},q_{p}$, combining its warm-phase component $q_{wc}$ and cold-phase component $q_{cc}$.
+Specifically, for vapor saturation, $q_{wc}=q_{vw}$ (saturation over water) and $q_{cc}=q_{vi}$ (saturation over ice); for cloud condensate, $q_{wc}=q_{w}$ (cloud water) and $q_{cc}=q_{i}$ (cloud ice); and for precipitation, $q_{wc}=q_{r}$ (rain) and $q_{cc}=q_{s}$ (snow).
+Building on this framework, we parameterize diverse microphysical processes, unified by the temperature-dependent factor $\alpha$.
+The saturation vapor mixing ratio $q_{vs}$ combines contributions from vapor saturation over water and ice(Yau and Rogers, 1996):
+(7) $q_{vs}=\frac{380.16}{p}\left[\alpha\exp\left(\frac{17.67\cdot T}{T+243.50}\right)+(1-\alpha)\exp\left(\frac{24.46\cdot T}{T+272.62}\right)\right],$
+where $p$ represents the pressure, and $T$ denotes the temperature.
+The saturation vapor is modeled as an exponential distribution for both liquid water and ice.
+The condensation rate $C_{C}$ governs cloud formation from water vapor, integrating water and ice contributions (Dudhia, 1989), where $\beta_{w},\beta_{i}$ are phase-specific coefficients:
+(8) $C_{C}=\beta_{w}\cdot\left(\alpha\cdot q_{vs}-q_{v}\right)^{+}+\beta_{i}\cdot\left((1-\alpha)\cdot q_{vs}-q_{v}\right)^{+}.$
+The autoconversion rate $A_{C}$ describes cloud condensate aggregation into precipitation (Kessler, 1995; Lin et al., 1983):
+(9) $A_{C}=\beta_{r}\cdot(\alpha\cdot q_{c}-10^{-3})^{+}+\beta_{s}\cdot e^{0.025T}\cdot((1-\alpha)\cdot q_{c}-10^{-3})^{+}.$
+The accretion rate $G_{C}$ models the collection of cloud condensate by precipitation particles (Morrison et al., 2015):
+(10) $G_{C}=q_{c}\cdot q_{p}\cdot\left(\beta_{r}\cdot\alpha^{2}+\beta_{s}\cdot(1-\alpha)^{2}\right).$
+The diffusional growth rate $D_{P}$ describes the conversion between precipitation and vapor through deposition and evaporation processes (Klemp and Wilhelmson, 1978), and is modeled as follows:
+(11) $D_{P}=\alpha(1-\alpha)\frac{q_{vs}-q_{v}}{\hat{\rho}q_{vs}}\frac{\left(\hat{\rho}q_{p}\right)^{0.525}}{5.4\cdot 10^{5}+2.55\cdot 10^{8}\left(pq_{vs}\right)^{-1}},$
+where $\hat{\rho}$ is the density of humid air, calculated in the same way as in (Hädrich et al., 2020), and $p$ represents the atmospheric pressure, which will be described in Section 4.3.
+The terminal velocity $\boldsymbol{u}_{p}$ of precipitation during deposition is modeled as a weighted combination of rain $\boldsymbol{u}_{r}$ and snow $\boldsymbol{u}_{s}$ velocities.Since precipitation particles quickly reach a uniform terminal velocity, we use typical values of $\boldsymbol{u}_{r}=-10\,\mathrm{m/s}$ for rain and $\boldsymbol{u}_{s}=-2\,\mathrm{m/s}$ for snow:
+(12) $\boldsymbol{u}_{p}=\alpha\boldsymbol{u}_{r}+(1-\alpha)\boldsymbol{u}_{s},$
+4.2. Hydrometeor Electrification
+Building on the Reynolds-Brook theory of thunderstorm electrification (Latham and Miller, 1965), the fair weather electric field induces opposing charges on precipitation particles.
+As these particles collide and coalesce, their velocities and directions influence partial charge neutralization, leaving a residual net charge.
+This theory underpins the following mathematical model for electrification, with charge density $\rho$ defined as:
+(13) $\rho=k\alpha(1-\alpha)\left(q_{c}^{2}\boldsymbol{u}_{c}-q_{p}^{2}\boldsymbol{u}_{p}\right)e^{-10q_{v}},$
+where $k$ is a user-defined constant, $\boldsymbol{u}_{c}$ and $\boldsymbol{u}_{p}$ are their respective velocities of cloud and precipitation.
+The threshold electric field $E_{tr}$ required for lightning initiation depends on the altitude $\rho_{A}(h)$ as defined by (Marshall et al., 1995):
+(14) $E_{tr}=\pm 167\,\rho_{A}(h)\quad\text{where}\quad\rho_{A}(h)=1.208\exp\left(\frac{-h}{8.4}\right).$
+To analyze the electrodynamics in the atmosphere, we solve the Poisson equation, a fundamental equation derived from Gauss’s law in electrostatics.
+This equation describes the simplified spatial variation of the potential $V$ due to charge density $\rho$(Kim and Lin, 2007):
+(15) $\nabla^{2}V=-4\pi\rho.$
+Charge neutralization processes, occurring after a lightning discharge(Barthe and Pinty, 2007), are governed by the following relationship:
+(16) $\delta\rho=\begin{cases}\pm\left(|\rho|-\rho_{\text{excess}}\right),&\text{if }|\rho|>\rho_{\text{excess}},\\ 0,&\text{if }|\rho|\leq\rho_{\text{excess}},\end{cases}$
+where $\delta\rho$ represents the net charge change, and $\rho_{\text{excess}}$ denotes the threshold for excess charge density.
 The total neutralized charge density, accounting for the collective contributions of lightning growth points, is expressed as:
-(19) $\delta\rho_{\text{neut}}=\frac{1}{N}\left|(2\alpha-1)q_{c}q_{p}\right|\sum_{n}\delta\rho$
-Finally, a suppression factor ($S$) is introduced to modulate lightning activity frequency as thunderstorms dissipate.
-This factor ensures consistent lightning activity during the lifecycle of thunderstorms:
-(20) $S=\frac{1}{1+k\cdot|\delta\rho_{\text{neut}}|}$
+(17) $\delta\rho_{\text{neut}}=\frac{1}{N}\left|(2\alpha-1)q_{c}q_{p}\right|\sum_{n}\delta\rho.$
+Finally, a suppression factor $S$ is introduced to modulate lightning activity as thunderstorms dissipate:
+(18) $S=\frac{1}{1+k\cdot|\delta\rho_{\text{neut}}|}.$
 4.3. Atmospheric Background
@@
 The isentropic exponent $\gamma_{\text{th}}$ for the air-water mixture (Anderson, 1990) is calculated as a weighted average of the vapor-specific exponent $\gamma_{\text{vapor}}$ and the air-specific exponent $\gamma_{\text{air}}$, as shown below:
-(21) $\gamma_{\text{th}}=Y_{\text{vapor}}\gamma_{\text{vapor}}+\left(1-Y_{\text{vapor}}\right)\gamma_{\text{air}}$
+(19) $\gamma_{\text{th}}=Y_{\text{vapor}}\gamma_{\text{vapor}}+\left(1-Y_{\text{vapor}}\right)\gamma_{\text{air}},$
 where $Y_{\text{vapor}}$ represents the mass fraction of water vapor in the air.
@@
 The temperature at a given altitude $h$ is expressed as:
-(22) $T(h)=\begin{cases}T_{0}+\Gamma_{0}h,&0\leq h\leq h_{1}\\ T_{0}+\Gamma_{0}h_{1}+\Gamma_{1}(h-h_{1}),&h_{1}\leq h\end{cases}$
+(20) $T(h)=\begin{cases}T_{0}+\Gamma_{0}h,&0\leq h\leq h_{1},\\ T_{0}+\Gamma_{0}h_{1}+\Gamma_{1}(h-h_{1}),&h_{1}\leq h,\end{cases}$
 where $T_{0}$ is the base temperature at sea level, $\Gamma_{0}$ and $\Gamma_{1}$ represent the lapse rates in the lower and upper layers.
 The atmospheric pressure profile, $p(h)$, is derived from the hydrostatic equation (Houze Jr, 2014), considering the effect of gravity and the ideal gas law. It is given by:
-(23) $p(h)=p_{0}\left(1-\frac{\Gamma h}{T_{0}}\right)^{\frac{g}{RT_{0}}}$
+(21) $p(h)=p_{0}\left(1-\frac{\Gamma h}{T_{0}}\right)^{\frac{g}{RT_{0}}},$
 where $p_{0}$ is the pressure at sea level, $g$ is the acceleration due to gravity, and $R$ is the specific gas constant.
@@
 This is given by:
-(26) $T_{\mathrm{th}}(h)=T_{0}\left(\frac{p(h)}{p_{0}}\right)^{\frac{\gamma_{\mathrm{th}}-1}{\gamma_{\mathrm{th}}}}$
+(24) $T_{\mathrm{th}}(h)=T_{0}\left(\frac{p(h)}{p_{0}}\right)^{\frac{\gamma_{\mathrm{th}}-1}{\gamma_{\mathrm{th}}}}.$
 Finally, the buoyancy force, which drives the upward movement of thundercloud, is calculated based on Archimedes’ principle and Newton’s second law.
@@
 […]
-4.4. Electrodynamics
-The electrodynamic behavior of the atmosphere is governed by Maxwell’s equations(Ma et al., 1998), which define the relationships between the electric field $\mathbf{E}$, charge density $\rho$, and electric potential $V$.
-Gauss’s law describes how the divergence of $\mathbf{E}$ is proportional to $\rho$:
-(28) $\nabla\cdot\mathbf{E}=\frac{\rho}{\epsilon_{0}}$
-where $\epsilon_{0}$ is the permittivity of free space. The electric field is related to the potential $V$ through:
-(29) $\mathbf{E}=-\nabla V$
-indicating that $\mathbf{E}$ points in the direction of decreasing potential. Together, these equations describe the electrostatic field and potential in regions with charge density.
-To analyze the electric potential in the atmosphere, we solve the Poisson equation, a fundamental equation derived from Gauss’s law in electrostatics.
-This equation describes the simplified spatial variation of the potential $V$ due to charge density $\rho$(Kim and Lin, 2007):
-(30) $\nabla^{2}V=-4\pi\rho$
-For simulating lightning discharges, the Dielectric Breakdown Model (DBM) is employed(Niemeyer et al., 1984).
-The probability of a discharge at a specific point $i$ depends on the electric potential $V_{i}$, expressed as:
-(31) $p_{i}=\frac{\left(V_{i}\right)^{\eta}}{\sum_{j=1}^{n}\left(V_{j}\right)^{\eta}}$
-Here, $\eta$ is a parameter controlling the spatial concentration of the discharge(Kim and Lin, 2007).
-Larger values of $\eta$ bias the discharge toward regions of higher potential, while smaller values allow for more diffusive propagation.
 4.4. Fluid Dynamics
-The motion of atmospheric fluids is described by the Navier-Stokes equations, which represent the conservation of momentum in a fluid medium (Wendt, 2008).
-These equations capture the effects of inertial forces, pressure gradients, viscosity, and external forces:
-(32) $\frac{\partial(\rho\mathbf{u})}{\partial t}+\nabla\cdot(\rho\mathbf{u}\mathbf{u})=-\nabla p+\rho\left(\mu\nabla^{2}\mathbf{u}+\mathbf{b}+\mathbf{f}\right),\quad\nabla\cdot\mathbf{u}=0$
-Here, $\rho$ is the fluid density, $\mathbf{u}$ the velocity field, $p$ the pressure, $\mu$ the dynamic viscosity, $\mathbf{b}$ the body force (e.g., gravity), and $\mathbf{f}$ external forces.
-The incompressibility condition $\nabla\cdot\mathbf{u}=0$ ensures mass conservation for incompressible flows.
+The motion of atmospheric fluids is governed by the incompressible Navier-Stokes equations, which represent the principles of momentum and mass conservation in a fluid medium (Wendt, 2008).
+These equations account for the effects of inertial forces, pressure gradients, viscosity, and external forces:
+(26) $\frac{\partial(\varrho\boldsymbol{u})}{\partial t}+\nabla\cdot(\varrho\boldsymbol{u}\boldsymbol{u})=-\nabla p+\varrho\left(\mu\nabla^{2}\boldsymbol{u}+\boldsymbol{f}\right),\quad\nabla\cdot\boldsymbol{u}=0,$
+where $\varrho$ represents the fluid density, $\boldsymbol{u}$ is the velocity, $p$ is the pressure, $\mu$ denotes the dynamic viscosity, and $\mathbf{f}$ encompasses buoyancy and other external forces.
+The equation $\nabla\cdot\boldsymbol{u}=0$ represents the continuity equation, ensuring the conservation of mass in incompressible flows.The pressure term, $p$, is determined via the Helmholtz–Hodge decomposition by solving the Poisson equation:
+(27) $\nabla^{2}p=\nabla\cdot\boldsymbol{u}.$
 5. ALGORITHMICS
-Algorithm 1 Thunderstorm Development Algorithm.
-1:Input: Current MCS state $(\theta,p,\rho,u,q_{v},q_{c},q_{p})$.
+The theoretical framework discussed in the previous section is translated into a numerical procedure, as outlined in Algorithm 1.
+Figure 2 visualizes the key interrelationships governing MCS development.
+We first present our algorithm for simulating the life cycle of a MCS, followed by an explanation of the implementation details of our procedure.
+5.1. Mesoscale Convective System Cycle
+The evolution of the MCS system follows Algorithm 1, iteratively updating state variables, including potential temperature $\theta$, pressure $p$, charge density $\rho$, velocity field $\boldsymbol{u}$, and hydrometeor quantities $q_{v}$, $q_{c}$, and $q_{p}$.
+The process begins with updating atmospheric background conditions $\theta,p,\rho$ following Eqs.
+(20–21).
+The velocity field $\boldsymbol{u}$ is then advected and diffused, followed by the computation of thermal buoyancy $b$ based on Eq.
+(25).
+Buoyancy, wind, and vorticity confinement forces are subsequently integrated into velocity field $\boldsymbol{u}$ to reflect key atmospheric dynamics.
+Algorithm 1 Thunderscapes Algorithm
+1:Input: Current MCS state $(\theta,p,\rho,\boldsymbol{u},q_{v},q_{c},q_{p})$.
 2:Output: Updated MCS state.
 3:procedure
-4:     $\theta,p,\rho\leftarrow$ Update atmospheric background conditions Eqs.(22–23)
-5:     $u\leftarrow$ Advect and diffuse velocity field
-6:     $b\leftarrow$ Compute thermal buoyancy Eq.(27)
-7:     $u\leftarrow u+b+f_{w}+f_{v}$triangleright Apply buoyancy, wind, vorticity confinement
+4:     $\theta,p,\rho\leftarrow$ Update atmospheric background conditions Eqs.(20–21)
+5:     $\boldsymbol{u}\leftarrow$ Advect and diffuse velocity field
+6:     $\boldsymbol{b}\leftarrow$ Compute thermal buoyancy Eq.(25)
+7:     $\boldsymbol{u}\leftarrow\boldsymbol{u}+\boldsymbol{b}+\boldsymbol{f}_{w}+\boldsymbol{f}_{v}$triangleright Apply buoyancy, wind, vorticity confinement
 8:     $q_{v},q_{c},q_{p}\leftarrow$ Advect hydrometeor quantities
-9:     $u\leftarrow$ Pressure projection
-10:     $q_{v},q_{c},q_{p},\theta,\rho\leftarrow$ Thundercloud microphysics update Eqs.(1–15)
-11:     if $S\cdot\rho>E_{tr}$ then $\triangleright$ Electrification process Eq.(16–18)
-12:         $V,\rho_{\text{neut}}\leftarrow$ lightning discharge Eqs.(30–31),(19)
-13:         $S\cdot\rho\leftarrow$ Apply lightning neutralization Eq.(20)
-14:     end if
-15:end procedure
-The theoretical framework discussed in the previous section is translated into a numerical procedure, as outlined in Algorithm 1.
-Figure 3 visualizes the key interrelationships governing thunderstorm development.
-We first present our methodology for simulating the life cycle of a Mesoscale Convective System (MCS), followed by an explanation of the dynamics of lightning channel growth triggered when the electric field exceeds a defined threshold.
-5.1. Mesoscale Convective System Cycle
-The input fields for the MCS system, depicted in Figure 5, serve as the foundation for initializing state variables, including potential temperature ($\theta$), pressure ($p$), charge density ($\rho$), velocity field ($u$), and hydrometeor quantities ($q_{v}$, $q_{c}$, $q_{p}$).The solver updates these variables iteratively, producing an output that represents the evolved state of the MCS, incorporating changes driven by thundercloud microphysics and the electrification processes.
-Figure 5.
-Different types of sources used as input for our MCS solver,temperature and vapor field are sampled from perlin noise patterns.
-(a) Height field derived from satellite data, (b) ground temperature distribution, (c) vapor field distribution.
-The procedure begins by updating the atmospheric background conditions,through governing equations (Eqs. 22–23).
-The velocity field is then advected and diffused.Thermal buoyancy is computed (Eq. 27) and integrated into the velocity field, alongside contributions from wind and vorticity confinement forces.
-Hydrometeor quantities ($q_{v}$, $q_{c}$, $q_{p}$) are transported according to the evolving velocity field through advection.
-Hydrometeor quantities ($q_{v}$, $q_{c}$, $q_{p}$) are transported through advection based on the evolving velocity field.
-For efficient pressure projection, we employ the compact Poisson filter scheme (Rabbani et al., 2022), a GPU-friendly method for solving large-scale sparse linear systems.
-Cloud microphysics are subsequently updated, incorporating processes such as condensation, evaporation, and precipitation, governed by parameterized equations (Eqs.
-1–15).
-Electrification is initiated when the charge density, modulated by the suppress factor ($S\cdot\rho$), exceeds the electric field threshold ($E_{tr}$, Eqs.
-16–18).
-This triggers lightning discharges, which are detailed in the next subsection.
-Following lightning activity, the neutralized charge density ($\rho_{\text{neut}}$, Eqs.
-30–31, 19) is updated, redistributing charge density ($S\cdot\rho$, Eq.
-20).
-The suppress factor is recalibrated to reflect the diminished likelihood of subsequent lightning events as the system dissipates.
-5.2. Lightning Channel Growth
-The growth of the lightning channel begins at a trigger point, randomly selected from locations where the electric field exceeds a predefined threshold.
-Inspired by the DBM model proposed by (Kim and Lin, 2007), the lightning channel stops growing once it connects to a positively charged object.
-This approach not only enhances flexibility in controlling the final structure of the lightning channel but also improves computational efficiency.
-The positive charge is distributed along the top of the height field.
-The solution is computed on a volumetric grid, and the results are subsequently converted from grid points to geometric points.
-To create a more natural branching structure, a random dithering technique is applied to these points.
-The points are then connected to form the final geometry of the lightning branches.
+9:     $\boldsymbol{u}\leftarrow$ Pressure projection Eq.(27)
+10:     $q_{v},q_{c},q_{p},\theta\leftarrow$ Update cloud microphysics and temperature Eqs.(2–5)
+11:     $\rho\leftarrow$ Hydrometeor electrification Eqs.(13)
+12:     if $S\cdot\rho>E_{tr}$ then Eq.(14)
+13:         $V,\rho_{\text{neut}}\leftarrow$ lightning discharge Eqs.(15–17)
+14:         $S\leftarrow$ Apply lightning neutralization Eq.(18)
+15:     end if
+16:end procedure
+Hydrometeor quantities $q_{v},q_{c},q_{p}$ are transported through advection in accordance with the updated velocity field $\boldsymbol{u}$.
+To ensure mass conservation, the velocity field $\boldsymbol{u}$ is corrected via pressure projection following Eq.
+(27).
+Next, cloud microphysics are updated by solving parameterized equations based on Eqs.
+(2–5) to account for condensation, evaporation, and precipitation.
+Hydrometeor electrification is modeled using parameterized equations following Eq.
+(13).
+When the charge density, modulated by the suppress factor $S\cdot\rho$, exceeds the electric field threshold $E_{tr}$ based on Eq.
+(14), lightning discharge is triggered.
+During the discharge, lightning channel growth is stochastically modeled using principles from the Dielectric Breakdown Model (DBM) (Kim and Lin, 2007).
+Finally, neutralized charge density $\rho_{\text{neut}}$ is updated following Eqs.
+(15–17), and the suppress factor is recalibrated to reduce the likelihood of subsequent lightning events as thunderstorms dissipate based on Eq.
+(18).
+5.2. Implementation
+Our tool is implemented as a Houdini Digital Asset (HDA) using Houdini 20.0.625’s microsolver framework, with OpenCL for GPU acceleration.
+The hardware setup comprises an NVIDIA® GeForce® RTX A6000 GPU, a 13th Gen Intel® Core™ i9-13900 processor, and 128GB of RAM.
+The framework is designed to align with the traditional workflow of digital artists, utilizing custom volume fields as input sources, including temperature field, vapor field, and height field.These fields, representing ground properties, serve as the foundation for driving the MCS simulation.
+Our framework operates on a uniform 3D grid, utilizing a semi-Lagrangian scheme to advect system quantities.
+To simulate horizontal thunderstorm movement, we apply a uniform horizontal wind field.
+Combined with vertical buoyancy forces, this approach produces a more realistic depiction of thunderstorm surges.
+To mitigate numerical dissipation, which can cause small-scale vortices to vanish prematurely, we integrate vorticity confinement techniques (Miyazaki et al., 2002).
+This enhancement ensures that fine-scale details, such as swirling motions within the storm, are preserved throughout the simulation.
+Lightning channel growth during hydrometeor discharge events is modeled using the Dielectric Breakdown Model (DBM) proposed by Kim et al.
+(Kim and Lin, 2007).
+The resulting lightning channels are converted from grid points to geometric points.
+To achieve a more natural branching structure, a random dithering technique is applied to these points, which are subsequently connected to form the final geometry of the lightning branches.
+Both the sparse Poisson problem for fluid pressure projection and the DBM processes are efficiently handled using the compact Poisson filter (Rabbani et al., 2022), a GPU-friendly method specifically optimized for solving large sparse linear systems.
 6. VISUALIZATION
-Our tool is implemented as a HDA(Houdini Digital Asset) using Houdini 20.0.625’s microsolver framework with OpenCL for GPU acceleration.
-The hardware setup includes NVIDIA® GeForce® RTX A6000 GPU,13th Gen Intel® Core™ i9-13900 processor, and 128GB of RAM.
-The simulation results demonstrate various thunderstorm and lightning phenomena, integrated with scenarios inspired by real-world weather events.
-As shown in Table 1, we summarize the lightweight parameter set used for the scenes presented in this section.
-All renders are produced using Houdini’s native volumetric rendering engine to ensure high-quality visual outputs.
+The simulation results demonstrate various thunderstorm phenomena, integrated with scenarios inspired by real-world weather events.
+As shown in Table 1, we summarize the spatial scale used for the scenes presented in this section.
+All renders are produced using Houdini’s native volumetric rendering engine to ensure high-quality visual outputs.For dynamic details, please refer to our supplementary video.
 Table 1.
-Overview of the lightweight parameter set used in the scenes presented in this section.
-$D$: World dimensions (km); $R$: Volume resolution; $T_{\text{mcs}}$: Computation time in seconds per frame(s).
+The table provides an overview of the spatial scale used in the scenes presented in this paper.
+Moreover, resolution ($R$) and computation time ($T_{\text{mcs}}$) measured in seconds per frame are listed.
+A constant time step size of $\Delta t=1\,\mathrm{min}$ is used.
+The background parameters defining the atmosphere are set to $\Gamma=-8.5\,\mathrm{K}/\mathrm{km}$ and $h_{1}=8\,\mathrm{km}$.
 Fig. Scene $D$ (km) $R$ $T_{mcs}$ (s)
-6(a) Single Cell $18\times 12\times 18$ $450\times 301\times 450$ 0.48
-6(b) Multicell $20\times 15\times 20$ $500\times 376\times 500$ 0.81
-6(c) Squall Line $30\times 15\times 30$ $500\times 251\times 500$ 0.54
-6(d) Super Cell $30\times 15\times 30$ $500\times 251\times 500$ 0.68
-7(a) Florida $11\times 12\times 11$ $287\times 309\times 287$ 0.24
-7(b) New Mexico $11\times 13\times 11$ $383\times 438\times 383$ 0.59
-7(c) Japan $11\times 13\times 11$ $383\times 444\times 383$ 0.64
-7(d) California $10\times 7\times 10$ $340\times 247\times 340$ 0.28
-6.1. Thunderstorms Variation
-We simulate four common types of thunderstorms that occur within a mesoscale convective system (MCS), capturing their distinctive characteristics and dynamic behaviors:
+ 3 Single Cell $18\times 12\times 18$ $450\times 301\times 450$ 0.48
+ 3 Multicell $20\times 15\times 20$ $500\times 376\times 500$ 0.81
+ 3 Squall Line $30\times 15\times 30$ $500\times 251\times 500$ 0.54
+ 3 Super Cell $30\times 15\times 30$ $500\times 251\times 500$ 0.68
+ 4 Florida $11\times 12\times 11$ $287\times 309\times 287$ 0.24
+ 4 New Mexico $11\times 13\times 11$ $383\times 438\times 383$ 0.59
+ 4 Japan $11\times 13\times 11$ $383\times 444\times 383$ 0.64
+ 4 California $10\times 7\times 10$ $340\times 247\times 340$ 0.28
+6.1. Thunderstorm Variation
+We simulate four common types of thunderstorms that occur within a MCS according to the National Severe Storms Laboratory (NSSL) [^1] https://www.nssl.noaa.gov/education/svrwx101/thunderstorms/types/, capturing their distinctive characteristics and dynamic behaviors:
   • Single Cell: Single cell thunderstorms, often referred to as “popcorn” convection, are small, isolated storms characterized by their brief lifespan, typically lasting less than an hour.
@@
 Refer to caption(h)
+Refer to caption(i)
+Refer to caption(j)
+Refer to caption(k)
+Refer to caption(l)
 Figure 3.
@@
 This layout highlights the structural similarities between our simulated storms and their real-world counterparts.
-These simulated variations allow us to explore the diverse behaviors and impacts of thunderstorms within an MCS.
-The visual comparison of these thunderstorm types is presented in Figure 6, highlighting their structural differences and unique features.
+Additionally, the last row (i-l) illustrates the ground properties (including the temperature field and vapor field initialized with a consistent pattern based on a flat heightfield) associated with each thunderstorm type.
+These simulated variations allow us to explore the diverse behaviors and impacts of thunderstorms within a MCS.
+Figure 3 presents a visual comparison along with the corresponding parameter sets, showcasing the structural differences and distinctive features of these thunderstorm types.
 6.2. Severe Weather Phenomena
-To explore the impact of thunderstorms in real-world scenarios, we reference severe weather events in geographically diverse locations.
-These events illustrate the variability and intensity of thunderstorms in different environments:
+To explore the impact of MCSs in real-world scenarios, we reference severe weather events from geographically diverse locations using the Storm Events Database provided by the National Oceanic and Atmospheric Administration (NOAA) [^2] https://www.ncdc.noaa.gov/stormevents.
+These events highlight the variability and intensity of thunderstorms across different environmental conditions and seasons:
+  • Biscayne National Park, Florida (25.3692°N, -80.243°W): On March 13, 1993, a tropical storm occurred in this region during the spring, bringing heavy rainfall and strong winds.
+  • Chiricahua Mountains, New Mexico (35.7943°N, -106.443°W): On July 14, 1989, a severe thunderstorm swept through this mountainous area during the summer, producing heavy rain, hail, and strong winds.
+  • Fuji-Hakone-Izu National Park, Japan (37.745°N, -119.54°W): On January 10, 2011, a sudden winter thunderstorm formed in this region, resulting in brief but intense snowfall.
+  • Monterey Bay, California (36.6197°N, -121.906°W): On December 2, 2012, an atmospheric river brought heavy rainfall and thunderstorms along the California coast, accompanied by strong gusty winds.
 Refer to caption(a)
@@
 Refer to caption(d)
+Refer to caption(e)
+Refer to caption(f)
+Refer to caption(g)
+Refer to caption(h)
 Figure 4.
-Simulation cloudscapes inspired by real-world weather events.
-(a) Florida Biscayne National Park, (b) New Mexico Chiricahua Mountains, (c) Japan Fuji-Hakone-Izu National Park, (d) California coastal atmospheric river.
-These visualizations highlight the geographical diversity and meteorological phenomena captured in our simulation framework.
-  • Biscayne National Park, Florida (25.3692, -80.243): On March 13, 1993, a tropical storm occurred in this region during the spring, bringing heavy rainfall and strong winds.
-  • Chiricahua Mountains, New Mexico (35.7943, -106.443): On July 14, 1989, a severe thunderstorm swept through this mountainous area during the summer, producing heavy rain, hail, and strong winds.
-  • Fuji-Hakone-Izu National Park, Japan (37.745, -119.54): On January 10, 2011, a sudden winter ;p‘thunderstorm formed in this region, resulting in brief but intense snowfall.
-  • Monterey Bay, California (36.6197, -121.906): On December 2, 2012, an atmospheric river brought heavy rainfall and thunderstorms along the California coast, accompanied by strong gusty winds.
+Simulation inspired by real-world weather events:(a) Florida’s Biscayne National Park, (b) New Mexico’s Chiricahua Mountains, (c) Japan’s Fuji-Hakone-Izu National Park, (d) California’s Monterey Bay.
+These visualizations highlight the geographical diversity and meteorological phenomena captured in our simulation framework.Additionally, the last row (e-h) illustrates the ground properties (including the temperature field and vapor field, both initialized with a consistent pattern derived from a specific regional heightfield) associated with each region.
 These events serve as the basis for our simulations, showcasing the capability of our model to replicate the diverse and complex phenomena associated with severe weather.
-The corresponding visualizations are presented in Figure 7.
+The corresponding visualizations and parameter sets are presented in Figure 4.
 7. VALIDATION
-We adopt the evaluation methods outlined in (Shen et al., 2020) and (Cesana et al., 2019), utilizing altitude-based quantitative data to analyze cloud behavior across different atmospheric layers.
-The results are presented in Figure 8, which illustrates the relationship between cloud fraction and storm height.
-Notably, the maximum cloud fraction corresponds to the storm’s peak altitude, consistent with the classical vertical distribution of cumulonimbus clouds in meteorology.
+The validation of our model involves presenting spatially simulated thunderstorm structures alongside meteorological characteristics and comparing the temporally simulated results with real-world weather data.
+This includes analyzing cloud fraction profiles to assess cloud formation and structure, tracking the evolution of cloud coverage against real-life data from national weather services [^3] https://www.visualcrossing.com/weather/weather-data-services/, and evaluating the temporal variation in lightning flash rates.
 Figure 5.
-Comparative visualization of cloud fractions under different storm systems.
-From left to right: single-cell, multi-cell, squall line, and supercell.
-Each panel highlights the cloud fraction structure typical of these systems.
-For further validation, we specifically selected four regions frequently studied in thunderstorm meteorology (Uman, 2001; Harris and Carvalho, 2018).
-Following the methodology from (Herrera et al., 2021), we examine the temporal evolution of the thunderstorm described in section 6.2.
-The evolution of cloud coverage is compared with real-world weather data from public sources, showing dynamic changes in the distribution of cloud cover.
-This comparison helps assess the accuracy of the model in capturing cloud coverage patterns observed in nature.we present the maximum storm development height along with the corresponding cloud fraction, providing a detailed overview of the storm’s vertical structure.
-This information is used to validate the model’s ability to simulate the typical altitude and distribution of clouds during storm development.
-In line with the evaluation method in (Formenton et al., 2013), we analyze the evolution of lightning flash rates to capture the electrical activity of thunderstorms.
-The flash rate increases as the storm develops and gradually declines to zero as the storm dissipates.
-This trend reflects the typical lifecycle of a thunderstorm and its charge dynamics.
-Additionally, the frequency of lightning occurrences is examined in relation to altitude.
-Figure 9 provides a comparison across Florida, New Mexico, Japan, and California, illustrating the model’s ability to capture a wide range of climatic behaviors and atmospheric dynamics, and confirming its consistency with observed storm structures and lightning patterns.
+Visualization comparing cloud fraction models and simulation results for four storm types: single cell, multi-cell, squall line, and supercell.
+Top row: Cloud fraction models, where orange represents our method, and blue represents the method of Hädrich et al.
+(Hädrich et al., 2020).
+Middle row: Results generated using Hädrich et al.
+(Hädrich et al., 2020).
+Bottom row: Results generated by our method.
+The visualization highlights differences in cloud fraction structures and thunderstorm characteristics captured by the two approaches.
+7.1. Thunderstorm structure
+We adopt the evaluation methods outlined in (Shen et al., 2020) and (Cesana et al., 2019), utilizing altitude-based quantitative data to analyze cloud behavior across various atmospheric layers.
+To ensure a fair comparison, we evaluate our framework alongside the cumulonimbus simulation model Stormscapes (Hädrich et al., 2020) using identical input fields and consistent initial conditions.
+The results, shown in Figure 5, illustrate the relationship between cloud fraction and the structure of thunderstorms at their maximum development height as modeled by the two approaches.
+While both models perform well for simpler thunderstorm types, such as single cell thunderstorms, notable differences arise for more complex systems, including multi-cell thunderstorms, squall lines, and supercells.
+At higher altitudes, the Stormscapes struggles to capture the expected extension of cloud coverage, particularly the anvil cloud, which is a critical feature of complex thunderstorms.
+The anvil, characterized by its solid appearance, sharp, well-defined edges, and downwind spread influenced by upper-level steering winds, is not accurately represented in Stormscapes.
+This limitation arises because the Stormscapes relies on the classical Kessler-style precipitation parameterization, which only accounts for phase transitions among vapor, rain, and cloud water.
+In contrast, our framework employs a Grabowski-style microphysics scheme, which incorporates a finer classification of phase transitions, including vapor, rain, snow, cloud water, and ice.
+This enhanced granularity enables more accurate modeling of atmospheric buoyancy and better aligns with meteorological observations, making our model particularly well-suited for realistic MCS simulations.Moreover, the Stormscapes model does not account for hydrometeor electrification processes, limiting its ability to simulate lightning dynamics during thunderstorm development and to reproduce consistent atmospheric phenomena.
+Figure 6.
+Comparison of MCS across four distinct regions: Florida, New Mexico, Japan, and California.
+Top row: Cloud coverage evolution over a 24-hour period during thunderstorm events.
+Middle row: Cloud fraction analysis, illustrating the structure and distribution of clouds at the maximum development height of thunderstorms.
+Bottom row: Lightning activity evolution, capturing the variations in lightning activity as thunderstorms develop.
+The figures highlight the unique patterns shaped by regional environmental conditions.
+7.2. Consistent Thunderstorm System
+We proceed to analyze and quantify the simulation data discussed in Section 6.2.Figure 6 provides a comparative overview of MCS characteristics across four regions: Florida, New Mexico, Japan, and California.
+Additionally, Figure 7 illustrates the development of MCS over specific evolution times, offering insights into temporal dynamics.
+Our analysis begins by comparing the simulated cloud coverage attributes with 24-hour observational meteorological data.
+Next, we assess the cloud fraction at the maximum development height of thunderstorms in each region, providing critical insights into the vertical structure and dynamics of the thunderstorm systems.
+Using the evaluation methodology described in (Formenton et al., 2013), we further analyze the evolution of lightning flash rates to capture the electrical activity within the thunderstorms.
+The flash rate follows a characteristic trend: increasing during the thunderstorm’s intensification phase, peaking at maturity, and gradually declining as the system dissipates.
+Additionally, we examine the altitude of lightning generation, revealing the vertical distribution of electrical activity.
+The alignment of lightning flash rates with the spatial and temporal evolution of the thunderstorm systems demonstrates the model’s consistency in realistically simulating the life cycle of MCS.
+Figure 7.
+Different regions’ MCS evolution over time.
+First row: Florida at times $t_{1}=6.56\text{h},t_{2}=9.58\text{h},t_{3}=13.44\text{h},t_{4}=20.51\text{h}$.
+Second row: New Mexico at times $t_{1}=6.12\text{h},t_{2}=11.45\text{h},t_{3}=17.70\text{h},t_{4}=22.32\text{h}$.
+Third row: Japan at times $t_{1}=5.36\text{h},t_{2}=9.34\text{h},t_{3}=10.36\text{h},t_{4}=10.38\text{h}$.
+Fourth row: California at times $t_{1}=9.89\text{h},t_{2}=11.04\text{h},t_{3}=11.80\text{h},t_{4}=16.26\text{h}$.
 8. CONCLUSION AND FUTURE WORK
-We have developed a comprehensive framework for simulating atmospheric phenomena during thunderstorms, focusing on thundercloud development, dissipation, and lightning generation through cloud electrification processes.
-This framework integrates a Grabowski-style extended warm cloud microphysics scheme with hydrometeor electrification processes to ensure the consistent coupling of cloud formation and lightning dynamics.
-The model simulates key atmospheric quantities, including microphysical properties such as vapor, cloud water, ice, and precipitated rain and snow, as well as electrodynamic properties like static charge distribution and lightning flashes.
-Our framework reproduces various thundercloud types, including single-cell, multicell, squall line, supercell, and atmospheric river formations, and simulates lightning dynamics using the Dielectric Breakdown Model (DBM), capturing both CG and IC lightning.
-Validation against real-world weather data demonstrates the model’s accuracy in simulating cloud structure, temperature evolution, cloud coverage, relative humidity, and lightning flash rates, with results benchmarked against national weather services and observed lightning activity.
-Figure 9.
-Comparison of cloud coverage evolution (top row) ,cloud fraction (middle row) and lightning activity evolution(bottom row)across four distinct regions: Florida, New Mexico, Japan, and California.
-The figures showcase the distinct patterns of atmospheric dynamics influenced by regional environmental conditions.
+We have developed a physically based simulation framework for visually realistic MCS.
+This framework integrates a Grabowski-style extended warm cloud microphysics scheme with hydrometeor electrification processes to ensure the consistent coupling of thunderstorms formation and lightning dynamics.
+Our framework reproduces various thunderstorms types within a MCS.Validation against real-world weather data demonstrates the model’s consistency in simulating cloud structure, cloud coverage, and lightning flash rates, with results benchmarked against national weather services.
 Future work will focus on expanding the framework to incorporate advanced atmospheric concepts and support a broader range of thunderstorm and lightning phenomena.
@@
 Furthermore, new lightning types such as anvil crawlers, bolts from the blue, and sheet lightning will be introduced, requiring refined electric field and branching models.
-Finally, scalability and real-time performance will be improved by leveraging advanced numerical techniques and hardware optimizations, enabling more detailed and efficient simulations.
+Finally, scalability and performance will be enhanced by leveraging advanced numerical techniques and improving hardware utilization.
+Currently, our numerical method employs a uniform grid.Transitioning to an adaptive grid approach (Raateland et al., 2022) appears to be a promising direction for achieving higher performance.